### v1

# Sufficiently Regularized Nonnegative Quartic Polynomials are Sum-of-Squares

Abstract
Hilbert’s 17th problem famously established that not all nonnegative polynomials admit an SoS representation. Hilbert also identified a few special classes in which nonnegativity and SoS are equivalent, such as univariate polynomials, quadratic polynomials, and bivariate quartic polynomials. In this paper, we extend this equivalence to several new subclasses of multivariate quartically-regularized polynomials and characterise the NP-hardness boundary of these special-structure polynomials. Namely, we consider the global optimization of multivariate symmetric cubic polynomials regularised by weighted quartic powers of the Euclidean norm. These special-structure polynomials arise as iterative subproblems in high-order tensor methods for nonconvex optimisation problems. We consider shifting these polynomials by their global optimum so as to make them nonnegative, and show that for sufficiently high regularization and under mild assumptions, these polynomials admit a Sums-of-Squares (SoS) representation. We also identify several structured subclasses of quartically-regularized cubic polynomials for which global optimality of the model implies that its nonnegativity is certified by an SoS decomposition, for all values of the regularization parameter (such as quadratic-quartic polynomials and quartic polynomials containing a special cubic term that can be decomposed into a product of a quadratic norm and a linear form). We provide counterexamples using quartic separable norms that showcase the crucial role of the Euclidean norm in these representations. We illustrate how to use our SoS-based certificates for the Taylor subproblems inside high-order tensor methods for nonconvex optimization, with encouraging results.
1 Introduction
A foundational milestone in real algebraic geometry is Hilbert’s seminal 1888 paper [30], which provides a complete classification of the degrees and dimensions for which every nonnegative polynomial admits a representation as a sum of squares of polynomials. In this work [30], Hilbert showed that there exist nonnegative polynomials that cannot be expressed as sums of squares, although explicit examples of such polynomials were only constructed nearly eighty years later. The study of the gap between nonnegativity and sums of squares remains an active area of research to this day.
Motivated by new applications and growing interest in tensor methods and high-order optimization, there has been growing engagement within the optimization community in understanding when nonnegative polynomials can be represented as sums of squares (SoS). In particular, the minimization of regularized polynomials lies at the core of all regularization-based $p$th-order optimization methods ($p\geq 2$) and is fundamental to polynomial optimization. In these methods, we consider the unconstrained, potentially nonconvex, optimization problem $\min_{x\in\mathbb{R}^{n}}f(x)$, where $f:\mathbb{R}^{n}\to\mathbb{R}$ is $p$-times continuously differentiable and bounded below, for some $p\geq 2$. The adaptive $p$th-order regularization method (ARpp) is known to achieve optimal global and local convergence rates for sufficiently smooth objectives, in both convex and nonconvex settings [19, 16, 7, 49]. In the ARpp method, the subproblem is constructed using the $p$th-order Taylor expansion $T_{p,x_{k}}(s)$ of the objective at the iterate $x_{k}$, leading to the regularized polynomial,
$\min_{s\in\mathbb{R}^{n}}\;m_{p}(s):=T_{p,x_{k}}(s)+\frac{\sigma}{p+1}\|s\|^{p+1},$ (ARpp Subproblem)
where $\|\cdot\|$ denotes the Euclidean norm unless otherwise specified, while $\sigma>0$ is an adaptive regularization parameter. In [19, 9, 16, 15], it was proved that when $\sigma$ is sufficiently large – that is, when the model is adequately regularized – the step obtained from minimizing the subproblem is guaranteed to produce a decrease in the objective function value, a key ingredient for ensuring global convergence and rate of the ARpp method. In the nonconvex setting, Birgin et al [9] and Cartis et al [19, 16] have shown that the ARpp method requires at most $O\left(\epsilon_{g}^{-\frac{p+1}{p}}\right)$ evaluations of $f$ and its derivatives to obtain an iterate satisfying $\|\nabla f(x_{k})\|\leq\epsilon_{g}$, and that this bound is optimal within the class of smooth objectives with Lipschitz continuous $p$th order derivatives [15, 16]. These bounds further improve when $f$ is convex, as shown by Nesterov [49] and others. Moreover, ARpp methods achieve $p$th order local rates of convergence faster local convergence rates than first- and second-order methods [24, 71]. This demonstrates that increasing the model degree and accuracy order to $p>2$ leads to improved evaluation complexity and local rates compared to first- and second-order methods, providing strong motivation for the interest and use of higher-order methods in practice [20]. However, the question of how to efficiently solve the subproblems in ARpp methods in ways that exploit problem structure has only mainly been addressed in the convex case (see Nesterov [49, 53, 2] and follow-up works) until recently [74, 21] and very much remains an ongoing topic of research, especially in terms of understanding the level of computational difficulty involved.
It is well known that for polynomials of degree four or higher, deciding nonnegativity is NP-hard (see, for example, the NP-hardness of testing matrix copositivity [44]). Also, Nesterov has shown that maximizing a cubic polynomial over the unit sphere is also NP-hard [12, 39]. Although these results illustrate the general intractability of nonnegativity for high-degree polynomials, regularized polynomials form a more subtle and structured subclass. Interestingly, Parrilo et al. [60, Sec. 5.1] observed numerically that certain quartically regularized test functions appear to become SoS once the regularization parameter is chosen large enough. However, these observations were empirical; no theoretical proof, nor any explicit bound on the regularization magnitude, have been established in the literature. Recently, Ahmadi et al. [1] proved that a sufficiently regularized locally convex polynomial is SoS-convex, and the algorithmic framework is given in [75]. However, when (ARpp Subproblem) is not positive semidefinite, the presence or absence of the SoS property for the ensuing regularised polynomial model is an open question in the literature. The question of whether a regularized (nonnegative) even-degree symmetric polynomial necessarily becomes SoS for large regularization values remains essentially unexplored.
In the case of $p=3$, we let $f_{0}:=f(x_{k})\in\mathbb{R}$, $g:=\nabla f(x_{k})\in\mathbb{R}^{n}$, $H:=\nabla^{2}f(x_{k})\in\mathbb{R}^{n\times n}$ , $T:=\nabla^{3}f(x_{k})\in\mathbb{R}^{n\times n\times n}$. Then the (ARpp Subproblem) becomes a quartically-regularized cubic polynomial in $n$ variables,
$\min_{s\in\mathbb{R}^{n}}m_{3}(s)=f_{0}+g^{T}s+\frac{1}{2}H[s]^{2}+\frac{1}{6}\,T[s]^{3}+\frac{\sigma}{4}\,\|s\|^{4},$ (Quartically Regularized Cubic Model)
where $f_{0}$, $g$, $H$, and $T$ denote the constant, linear, quadratic, and cubic coefficients, respectively. Throughout, we assume that $H$ and $T$ are symmetric. Note that $H[s]^{2}$ (equivalently, $s^{T}Hs$) and $T[s]^{3}$ represent the second- and third-order directional derivatives at $x_{k}$ along $s\in\mathbb{R}^{n}$, respectively. The regularization term is a quartic-degree Euclidean norm, that is, $\frac{\sigma}{4}\|s\|^{4}=\frac{\sigma}{4}\,(s_{1}^{2}+\dotsb+s_{n}^{2})^{2}.$ Since $\sigma>0$, the model $m_{3}(s)$ is bounded below and the optimization problem is well defined. We denote the global minimum value of $m_{3}$ by $m_{*}$, and the global minimum is attained (see Remark 2.3). In this work, we aim to address the following fundamental question
When (Quartically Regularized Cubic Model) is sufficiently regularized (i.e., when $\sigma>0$ is large enough), is the shifted polynomial $m_{3}-m_{*}$ a sum of squares (SoS)?
Once a polynomial can be written as a sum of squares, this not only certifies its global nonnegativity but also allows the associated global polynomial optimization problem to be reformulated as a semidefinite program (SDP) at the zeroth level of the Lasserre hierarchy [33, 60]. Such SDPs have polynomial size $\binom{n+4}{4}$ and can be solved in polynomial time using, for example, interior-point methods [45]. Moreover, if the global minimizer is unique, it can be extracted in polynomial time through the moment–SoS hierarchy [36, 34].
In this work, we characterize the NP-hardness boundary of quartically regularized polynomials, thereby extending Hilbert’s classical characterizations to a new subfamily of multivariate polynomials. The contributions of the paper are as follows.
Regularization threshold for SoS: Under mild assumption, we establish an explicit bound on the regularization parameter $\sigma$ such that, if $\sigma$ exceeds this threshold, the shifted quartically regularized polynomial $(m_{3}-m_{*})$ is SoS. To the best of our knowledge, this is the first work to establish SoS results for a generally nonconvex quartically-regularized cubic polynomial; thus, building an understanding of how strength of regularization governs convexification, SoS certificates, and global optimality.
Special polynomial classes with equivalence between global optimality and SoS certificates: Hilbert’s 17th problem famously established that not all nonnegative polynomials admit an SoS representation [30]. Hilbert also identified a few special classes in which nonnegativity and SoS are equivalent, such as univariate polynomials, quadratic polynomials, and bivariate quartic polynomials. In this paper, we extend this equivalence to several new variants of multivariate quartically regularized polynomials. Specifically, (i) quadratic–quartic polynomials with any number of variables, and (ii) quartic polynomials containing a special cubic term that can be decomposed into a product of a quadratic norm and a linear form. These cases constitute a rare subclass of nonconvex polynomials that are SoS-certifiable, establishing the relationship between global optimality conditions and SoS representability.
Discussion of quartic norms: Lastly, we identify a subtle yet important distinction between different types of regularization norms. Several alternative regularization norms and seminorms have been proposed in the literature. For instance, Schnabel et al. [66] introduced a quartic seminorm formulation for unconstrained optimization of the form $\frac{1}{4}\sum_{j=1}^{k}\sigma_{j}(a_{j}^{T}s)^{4}$, where $\sigma_{j}>0$ and $a_{j}\in\mathbb{R}^{n}$. [76] discusses the separable quartic norm $\|s\|_{4}^{4}=\sum_{i=1}^{n}s_{i}^{4}$. We use explicit examples to illustrate that when the regularization term is not the Euclidean quartic norm $\|s\|^{4}$ but the separable quartic norm $\|s\|_{4}^{4}$, even for arbitrarily large $\sigma$, the corresponding nonnegative separable quartic polynomial need not be SoS. We provide detailed characterization and examples on the choice of regularization norm that leads to SoS representability.
The paper is organized as follows. Section 1.1 provides background and definitions on the SoS framework, the Lasserre hierarchy, and related results by Ahmadi, Lasserre, Nesterov, and Parrilo. Section 2 presents our main theoretical results, including explicit bounds on $\sigma$ under which the quartically regularized cubic polynomial becomes SoS. We further discuss how this bound differs between locally convex cases ($H\succ 0$) and locally nonconvex cases. Section 3 analyzes several special classes of quartically regularized polynomials, identifying which forms are SoS and which are not, and highlighting the distinctions arising from different regularization norms. Section 4 provides detailed examples and preliminary numerical illustrations illustrating how various regularization norms lead to different NP-hardness behaviors. Finally, Section 5 concludes the paper.
1.1 Preliminaries and Background on Nonnegative and SoS Polynomials
Let $\mathcal{P}_{n}^{\deg(h)}$ denote the class of real polynomials $h(s):\mathbb{R}^{n}\to\mathbb{R}$ with $n$ variables and degree $\deg(h)$ and let $h(s):\mathbb{R}^{n}\to\mathbb{R}$ be a (multivariate) real polynomial in variables $s=[s_{1},\dotsc,s_{n}]^{T}\in\mathbb{R}^{n}$. The degree of a polynomial, $\deg(h)$, is defined to be the highest degree of its component monomials [^1] The degree of a monomial $s^{\alpha}$ where $\alpha=[\alpha_{1},\dotsc,\alpha^{n}]$ is equal to $\alpha_{1}+\dotsc+\alpha_{n}$. If the global minimum exists and is attained, we denote it by $h^{*}:=\min_{s\in\mathbb{R}^{n}}h(s)$ and $s^{*}$ is a global minimizer. A homogeneous polynomial is a polynomial where all the monomials have the same degree. A polynomial $h(s)$ is said to be nonnegative if $h(s)\geq 0$ for all $s\in\mathbb{R}$.
Definition 1.1.
(SoS polynomial [1, 5, 32]) $h(s)$ is SoS if there exist polynomials $\tilde{h}_{1},\dotsc,\tilde{h}_{r}:\mathbb{R}^{n}\rightarrow\mathbb{R}$, for some $r\in\mathbb{N}$, such that $h(s)=\sum_{j=1}^{r}\tilde{h}_{j}(s)^{2}$ for all $s\in\mathbb{R}^{n}$. We refer to a problem as being SoS-exact at level 0 if $h(s)-h^{*}$ is a sum of squares polynomial.
Definition 1.1 implies that the degree of an SoS polynomial $h(s)$ has to be even, $\deg(h)=2d$, and that the maximum degree of each $\tilde{h}_{j}$ is $\frac{\deg(h)}{2}=d$. Let $\mathbb{R}[s]^{n\times n}$ be the real vector space of $n\times n$ real polynomial matrices, where each entry of such a matrix is a polynomial with real coefficients.
Definition 1.2.
(SoS-convex, [5, Def. 2.4]) $h(s)$ is SoS-convex if its Hessian $\nabla^{2}h(s)$ is an SoS-matrix. A symmetric polynomial matrix $\check{H}(s)\in\mathbb{R}[s]^{n\times n}$ is an SoS-matrix if there exists a polynomial matrix $\check{V}(s)\in\mathbb{R}[s]^{n_{1}\times n}$ for some $n_{1}\in\mathbb{N}$, such that $\check{H}(s)=\check{V}(s)^{T}\check{V}(s)$.
Clearly, if $h(s):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is SoS-convex, then $h(s)$ is convex [5, 32]. Recently, Ahmadi et al. [1] proved that for a locally strongly convex quartically regularized polynomial ( $H\succeq\delta I_{n}\succ 0$), there exists a sufficiently large regularization parameter $\sigma$ such that the polynomial becomes SoS-convex (see Theorem 1.1).
Theorem 1.1.
([1], Lemma 3 and Thm 3) If $H\succeq\delta I_{n}\succ 0$, then the following sum of squares programming
[…]
is feasible. Moreover, this program can be reformulated as an SDP of size that is polynomial in $n$.
By choosing $\sigma$ under this setup, the resulting polynomial is strongly convex, its global minimizer exists and is unique. Using the SoS-convexity, we deduce that the unique global minimizer can be obtained by solving an SDP whose size is polynomial in $N=\binom{n+2d}{2d}$. Therefore, for locally convex regularized polynomials, both determining such a regularization parameter $\sigma$ and globally minimizing the resulting SoS-convex polynomial reduce to solving semidefinite programs in polynomial time. The following lemma from [28] connects nonnegative SoS-convex polynomials to SoS polynomials, and therefore ensures that a sufficiently regularized, nonnegative, locally strongly convex polynomial is SoS.
Lemma 1.1.
(SoS Convex + Non-negativity $\Rightarrow$ SoS, Lemma 8 in [28]) Assume the global minimizer $s^{*}$ of $h(s)$ exists and $\nabla h(s^{*})=0$. Let $q(s):=h(s)-h(s^{*})$. Then, clearly, $q(s):=h(s)-\nabla h(s^{*})$ is a non-negative polynomial with $q(s^{*})=0$ and $\nabla q(s^{*})=0$. Moreover, if $q(s)$ is SoS convex (i.e., $\nabla^{2}h(s)$ is an SoS-matrix), then $q(s)$ is SoS.
While convexity often leads to tractable global optimization [59, 5, 49], local nonconvexity makes the problem markedly more difficult. In 1992, the question of determining whether a polynomial defines a convex function was posed as one of the seven major open problems in the complexity theory of numerical optimization [58]. Ahmadi, Olshevsky, and Tsitsiklis later proved that this decision problem is strongly NP-hard even for quartic polynomials [4]. A natural and fundamental question, first posed and (partially) answered by Hilbert, is to determine for which dimensions $n$ and degrees $\deg(h)$, nonnegative polynomials or forms can always be expressed as sums of squares.
Theorem 1.2 (Hilbert, [30]).
For the space of real polynomials $\mathcal{P}_{n}^{\deg(h)}$, nonnegativity and sum-of-squares coincide if and only if one of the following holds:
  1. 1.
    $n=1$ (univariate polynomials of any degree);
  2. 2.
    $\deg(h)=2$ (quadratic polynomials in any number of variables);
  3. 3.
    $n=2$ and $\deg(h)=4$ (bivariate quartic polynomials).
For all other combinations of dimension and degree, Hilbert showed that nonnegativity does not imply a sum-of-squares representation. However, within the class of quartic polynomials, the SoS behavior of quartically regularized models is largely unexplored. Explicit examples of nonnegative but non-SoS polynomials appeared in the 1960s, most notably the Motzkin polynomial $M(s_{1},s_{2})=s_{1}^{4}s_{2}^{2}+s_{1}^{2}s_{2}^{4}-3s_{1}^{2}s_{2}^{2}s_{3}^{2}+s_{3}^{6},$ a nonnegative trivariate polynomial of degree six [43]. A more recent example appears in  [2, Thm. 3.3], a nonnegative trivariate quartic polynomial which is not SoS,
$\tilde{m}_{\mathcal{A}}(s)=\frac{9}{4}+2\|s\|^{2}+8(s_{1}s_{2}+s_{1}s_{3}+s_{2}s_{3})+(s_{1}^{4}+s_{2}^{4}+s_{3}^{4}).$ (1)
Using this example and related constructions, Ahmadi [2] proved that testing nonnegativity of separable quartic–quadratic polynomials is NP-hard. Note the distinction between (1) and (Quartically Regularized Cubic Model): the former uses a separable quartic norm regularization, whereas the latter employs a Euclidean quartic norm. This leads to subtle but important differences in the characterization of their SoS representations. Also, Nesterov [46] proved that even very structured nonconvex polynomial optimization problems are NP-hard to solve globally. In particular, he showed that the following problems are NP-hard.
Theorem 1.3.
([46, Thm 4], [39]) The maximization of homogeneous cubic or quartic form over the Euclidean ball, and the optimization of a quartic homogeneous polynomial under quadratic constraints are NP-hard [^2] The problem formulations are given in (52)–(54)..
A common feature among these NP-hard optimization problems in Footnote 2 is that they use constraints—such as Euclidean ball constraints, trust-region constraints, or quadratic constraints. In all three cases, intractability arises when the global minimizer is on the boundary of the feasible set, i.e., the constraint is active. In contrast, the regularization framework in (ARpp Subproblem) adopts a different globalization strategy by replacing explicit constraints with a regularization term. This guarantees boundedness from below and yields a well-defined unconstrained optimization problem. We show in later sections that, when the regularization parameter is sufficiently large, minimizing (Quartically Regularized Cubic Model) becomes tractable.
1.1.1 Connection to SDP and Matrix Generalizations
As mentioned earlier, the computational advantage of sums of squares is the connection to semidefinite programming. It is well known that a polynomial is a sum of squares if and only if it admits a positive semidefinite Gram matrix representation.
Theorem 1.4.
(Parrilo [62]) Let $h\in\mathcal{P}_{n}^{2d}$ and $\phi_{d}(s)$ be the vector of all monomials of degree at most $d$. A polynomial $h(s)$ is SoS if and only if there exists a symmetric matrix $\cal Q$} such that (i) $h(s)=\phi_{d}(s)^{T}{\cal Q}\phi_{d}(s)$ for all $s\in\mathbb{R}^{n}$, (ii) ${\cal Q}\succeq 0$. Note that $\cal Q$} has size $\binom{n+2d}{2d}$.
Example 1.1.
For quartic polynomials ($d=2$), $\phi_{2}$ is a list of monomial basis with highest degree $2$,
$\phi_{2}(v):=[1,v_{1},\dotsc,v_{n},v_{1}^{2},\dotsc,v_{n}^{2},v_{1}v_{2},\dotsc,v_{1}v_{n},v_{2}v_{3},\dotsc,v_{2}v_{n},\dotsc,v_{n-1}v_{n}]^{T}\in\mathbb{R}^{\frac{1}{2}(n+2)(n+1)}.$ (2)
In this paper, we use simplified notation. Let $\omega\in\mathbb{R}^{\frac{1}{2}(n+2)(n+1)}$
$\omega:=[1,v,u,z]^{T}=[1,\underbrace{v_{1},\dotsc,v_{n}}_{:=v\in\mathbb{R}^{n}},\underbrace{u_{1},\dotsc,u_{n}}_{:=u\in\mathbb{R}^{n}},\underbrace{z_{1,2},\dotsc,z_{n-1,n}}_{:=z\in\mathbb{R}^{\frac{n(n-1)}{2}}}]^{T}.$ (3)
where $v=[v_{1},\dotsc,v_{n}]^{T}$, $u=[v_{1}^{2},\dotsc,v_{n}^{2}]$ and $z=\{z_{ij}\}_{1\leq i<j\leq n}=[v_{1}v_{2},\dotsc,v_{1}v_{n},v_{2}v_{3},\dotsc,v_{2}v_{n},\dotsc,v_{n-1}v_{n}]$.
If the nonnegative polynomial $h(s)-h^{*}$ admits an SoS representation, then—by the Lasserre moment–SoS hierarchy [1, 35, 60]—two semidefinite programs coexist in duality. The SoS primal formulation [35] takes the form of a single convex semidefinite program (SDP) with dimension $N=\binom{n+2d}{2d}$.
$\min_{h^{*},\,{\cal Q}\succeq 0}\;h^{*}\quad\text{s.t.}\quad h(s)-h^{*}=\phi_{d}(s)^{T}{\cal Q}\,\phi_{d}(s),$ (4)
where $\phi_{d}(s)$ denotes the vector of all monomials in $s$ of degree $\leq d$, and ${\cal Q}\succeq 0$ is the associated Gram matrix. The global minimum $h^{*}$ is computable in polynomial time in the problem size. If, in addition, the global minimizer $s^{*}$ is unique, then the associated optimal moment matrix satisfies the rank condition, then we can recover the global minimizer $s^{*}$ in polynomial time [35, 60, 56] by solving the moment dual SDP formulation. As the resulting SDPs have size $N=\binom{n+2d}{2d}$, they could suffer from the curse of dimensionality, making SoS programming impractical for larger-scale problems ($n,d\geq 10$) without exploitable structure. When sparsity is present, sparsity-adapted SDP hierarchies can be used [13, 37, 41, 69]. Scalable variants such as DSoS/SDSoS and BSoS [3, 38, 69, 70, 72] further reduce SDP size by leveraging structural properties. Reformulation can also improve tractability—for example, a quartic homogeneous polynomial with quadratic constraints admits a quadratic SDP relaxation [39], involving only an $n\times n$ matrix instead of the $N\times N$ Gram matrix required in SoS.
Lastly, we remark that when the nonnegative polynomial $h(s)-h^{*}$ is not SoS, Lasserre [33] showed that a sequence of SDP relaxations can globally solve polynomial optimization problems, and Nie [57] further proved that Lasserre’s hierarchy has finite convergence. Thus, the nonconvex global optimization problem $\min_{s\in\mathbb{R}^{n}}h(s)$ can be approximated by a finite hierarchy of SDPs of increasing size. However, Scheiderer [65] proved that there is no uniform bound on the number of relaxation layers for finite convergence. As a consequence, optimizing a non-SoS polynomial $h(s)-h^{*}$ remains NP-hard.
1.1.2 Related Work on Local Methods and Iterative High-order Solvers
As we mentioned at the start of Section 1, there is a growing body of literature on high-order optimization methods, in both convex and nonconvex cases. We are particularly focused on reviewing here the works centered on developing iterative solvers and locally minimizing the regularized polynomials and subproblems involved in these methods. Nesterov and collaborators have made extensive contributions to convex higher-order optimization [49, 50, 51, 52, 53, 47], developing subproblem solution methods based on Bregman gradient techniques, higher-order proximal-point operators, and iterative minimization of convex quadratic models with quartic regularization (QQR). Recent developments by Cartis and Zhu introduced the nonconvex QQR method for minimizing nonconvex AR33 subproblems [22, 74]. Following QQR, the Cubic–Quartic Regularization (CQR) method [21] and the Diagonally Tensor Method (DTM) framework [76] were proposed, both leveraging cubic information to approximate local tensor structure. Local convergence guarantees for $p$th-order methods are analyzed in [24, 71], consistent with the local convergence theory of higher-order Newton methods in [1]. Tensor-based optimization techniques—including homogeneous polynomial formulations, generalized tensor eigenvalue problems, and tensor minimization tools—are investigated in [23, 63, 55, 54].
Recently, Ahmadi and Zhang [6] established polynomial-time checkable necessary and sufficient conditions for local minimality of cubic polynomials, implying that when a local minimizer exists, it can be found in polynomial time. Building on this, Silina et al. [68] proposed an unregularized third-order Newton method in which each iteration computes a local minimizer of the third-order Taylor model via an SDP in polynomial size. This unregularized approach coincides with the ARpp model when the regularization parameter is set to zero. A comprehensive overview of methods involving third- and higher-order derivatives is provided in [16, Ch. 4].
2 Nonnegative Quartically Regularized Polynomial is SoS for large $\sigma$
This section presents the main theoretical proof showing that the quartically regularized polynomial is SoS for sufficiently large $\sigma$. We begin by outlining the structure of the proof and the organisation of this section. First, we express the quartic polynomial
$q(v):=m_{3}(s^{*}+v)-m_{3}(s^{*})$
where $m_{3}$ is defined in (Quartically Regularized Cubic Model) via its fourth-order Taylor expansion at $s^{*}$ (see Lemma 2.1). We then establish sufficient conditions under which $m_{3}(s^{*}+v)-m_{3}(s^{*})$ admits an SoS representation (Section 2.1). These conditions require the usual first-order criticality condition $\nabla m_{3}(s^{*})=0$, together with a positive-semidefinite requirement $B(s^{*})\succeq 0$, where $B(s^{*})$ contains both the second-order term $\nabla^{2}m_{3}(s^{*})$ and the regularization contribution $\sigma\|s^{*}\|^{2}I_{n}$. The key idea is that, under mild assumption ($\|g\|\neq 0$), the regularization term $\sigma\|s^{*}\|^{2}I_{n}$ becomes dominant as $\sigma\to\infty$, eventually guaranteeing that $B(s^{*})\succeq 0$. To make this precise, in Section 2.2 we derive upper and lower bounds on $\|s^{*}\|$ and show that $\|s^{*}\|$ scales as $O(\sigma^{-1/3})$ as $\sigma\to\infty$, implying that $\sigma\|s^{*}\|^{2}\rightarrow\infty$ as $\sigma\to\infty.$ Finally, in Section 2.3 (see Theorem 2.3), when $\|g\|\neq 0$, we use these bounds to derive an explicit condition on $\sigma$ that ensures $B(s^{*})\succeq 0$, and consequently guarantees that the quartically regularized polynomial is SoS. In the case when $g=0$ (Remark 2.4), we are able to characterize several sub-cases in which $m_{3}(s)-m_{3}(s^{*})$ remains a sum of squares (though the general case of $m_{3}(s)-m_{3}(s^{*})$ with $g=0$ is dedicated to future work).
Theorem 2.1 and Theorem 2.3 are the two central results of this section. A detailed case-by-case analysis for the locally convex and locally nonconvex settings is provided in Section 2.3.1 and Section 2.3.2. These sections build the connections between SoS representations, convexification, SoS-convexity, and global optimality, and they characterize the resulting behaviour of the global minimizer in each regime.
2.1 SoS Representation of $m_{3}(s)-m_{*}$ and Characterizations of $s^{*}$
Using the global optimality conditions of $m_{3}$ in [76], we have the following useful result.
Lemma 2.1.
(Adapted from [76, Lemma 2.1]) Let $q(v):=m_{3}(s^{*}+v)-m_{3}(s^{*})$ where $m_{3}$ is defined in (Quartically Regularized Cubic Model). Then the fourth-order Taylor expansion of $q(v)$ at the point $s^{*}$ is exact and yields, for any vector $v\in\mathbb{R}^{n}$,
$q(v)$ $=\nabla m_{3}(s^{*})^{T}v+\frac{1}{2}\,\mathcal{G}(s^{*})[v]^{2}+\frac{1}{6}\,T[v]^{3}+\frac{\sigma}{4}\!\left(\|s^{*}+v\|^{2}-\|s^{*}\|^{2}\right)^{\!2}$
$=\nabla m_{3}(s^{*})^{T}v+\frac{1}{2}\,\nabla^{2}m_{3}(s^{*})[v]^{2}+\frac{1}{6}\,T[v]^{3}+\frac{\sigma}{4}\|v\|^{4}+\sigma\,{s^{*}}^{T}v\,\|v\|^{2},$ (5)
where $\mathcal{G}(s):=H+T[s]+\sigma\|s\|^{2}I_{n}$. Note that the first- and second-order derivatives of $m_{3}$ are
$\nabla m_{3}(s)=g+Hs+\frac{1}{2}\,T[s]^{2}+\sigma\|s\|^{2}s,\qquad\nabla^{2}m_{3}(s)=H+T[s]+\sigma\bigl(2ss^{T}+\|s\|^{2}I_{n}\bigr).$
where $I_{n}\in\mathbb{R}^{n\times n}$ is the identity matrix.
Remark 2.1.
Here we draw some connections with alternative formulations/expansions of $m_{3}$ and characterizations of $s^{*}$. In Lemma 8 in [28], the following result is shown: if $s^{*}$ is such that $\nabla m_{3}(s^{*})=0$, then
$m_{3}(s^{*}+v)-m_{3}(s^{*})=v^{T}\bigg[\underbrace{\int_{0}^{1}\int_{0}^{t}\nabla^{2}m_{3}(s^{*}+rv)}_{:=\mathcal{I}(s^{*},v)}drdt\bigg]v.$ (6)
Also, $s^{*}$ is the global minimum if and only if $\nabla m_{3}(s^{*})=0$ and $\mathcal{I}(s^{*},v)\succeq 0$ for all $v\in\mathbb{R}^{n}$.
The latter is equivalent to Lemma 2.1. We can see this using the following deductions. Using $\nabla^{2}m_{3}(s)=H+T[s]+\sigma(\|s\|^{2}+2ss^{T})$, from (6), we have
$\mathcal{I}(s^{*},v)$ $=$ $\int_{0}^{1}\int_{0}^{t}\bigg[H+T[s^{*}+rv]+\sigma\bigg(\|s^{*}+rv\|^{2}I_{n}+2[s^{*}+rv][s^{*}+rv]^{T}\bigg)\bigg]drdt$ (7)
$=$ $\int_{0}^{1}\int_{0}^{t}\bigg[\nabla^{2}m_{3}(s^{*})+r\bigg(T[v]+2\sigma{s^{*}}^{T}vI_{n}+4\sigma{s^{*}}v^{T}\bigg)+r^{2}\sigma(\|v\|^{2}I_{n}+2vv^{T}\big)\bigg]drdt$
$=$ $\frac{1}{2}\nabla^{2}m_{3}(s^{*})+\frac{1}{6}T[v]+\frac{\sigma}{6}\bigg(2{s^{*}}^{T}vI_{n}+4s^{*}v^{T}\bigg)+\frac{\sigma}{12}(\|v\|^{2}I_{n}+2vv^{T}\big).$
Multiplying $\mathcal{I}(s^{*},v)$ with $v^{T}$ and $v$ ,
$\mathcal{I}(s^{*},v)[v]^{2}$ $=$ $\frac{1}{2}\nabla^{2}m_{3}(s^{*})[v]^{2}+\frac{1}{6}T[v]^{3}+\sigma{s^{*}}^{T}v\|v\|^{2}+\frac{\sigma}{4}\|v\|^{4}$
which is the same result in (5).
Next, we present a technical lemma that rewrites the homogeneous cubic term $T[v]^{3}$, as a sum of linear polynomials multiplied by quadratic polynomials.
Lemma 2.2.
Any homogeneous cubic polynomial, $c_{3}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ with $\deg(c_{3})=3$ can be expressed as
$c_{3}(v)=\bigg(\sum_{i=1}^{n}u_{i}\bigg)\big(t^{T}v\big)+\bigg[\sum_{1\leq i<j\leq n}z_{ij}\big({v^{T}b^{(ij)}}\big)\bigg],$
where $u_{i}$ and $z_{ij}$ are quadratic terms, $t$ and $b^{(ij)}$ are vectors in $\mathbb{R}^{n}$ based on the coefficients of $c_{3}(v)$.
Proof.
For a homogeneous cubic polynomial, $c_{3}(v)=\frac{1}{6}T[v]^{3}$, we can write it in expanded form as $c_{3}(v):=\sum_{1\leq i\leq j\leq k\leq n}^{n}t_{ijk}v_{i}v_{j}v_{k}.$ Note that coefficients in the expanded form $\sum_{1\leq i\leq j\leq k\leq n}^{n}t_{ijk}v_{i}v_{j}v_{k}$ correspond to the tensor entries $\{T\}_{ijk}$ in tensor form $\frac{1}{6}T[v]^{3}$ as follows, $t_{ijk}=T_{ijk}$ for all distinct ijkijk, $t_{iij}=\frac{1}{2}T_{iij}$ and $t_{iii}=\frac{1}{6}T_{iii}$. We can further express $c_{3}(v)$ as a sum of linear polynomials times quadratic polynomials
$c_{3}(v):=\sum_{i,j,k=1,i\leq j\leq k}^{n}t_{ijk}v_{i}v_{j}v_{k}$ $=$ $\bigg(\sum_{i=1}^{n}v_{i}^{2}\bigg)\big(t^{T}v\big)+\bigg[\sum_{1\leq i<j\leq n}v_{i}v_{j}\big({v^{T}b^{(ij)}}\big)\bigg]$ (8)
$=$ $\bigg(\sum_{i=1}^{n}u_{i}\bigg)\big(t^{T}v\big)+\bigg[\sum_{1\leq i<j\leq n}z_{ij}\big({v^{T}b^{(ij)}}\big)\bigg]$
where $t=\{t_{iii}\}_{1\leq i\leq n}\in\mathbb{R}^{n}$. For $1\leq i<j\leq n$, $b^{(ij)}=[t_{1ij},\dotsc,t_{iij}^{*}\dotsc,t_{ijj}^{*}\dotsc,t_{ijn}]^{T}\in\mathbb{R}^{n}$ with $t_{iij}^{*}=t_{iij}-t_{jjj}$ and $t_{ijj}^{*}:=t_{ijj}-t_{iii}$ . ∎
Theorem 2.1 provides a sufficient condition under which $q(v):=m_{3}(s^{*}+v)-m_{3}(s^{*})$ can be expressed as a SoS polynomial, and it also gives an explicit SoS representation when the condition is satisfied.
Theorem 2.1.
Assume $\sigma>\nu>0$. Let $s^{*}$ satisfy $\nabla m_{3}(s^{*})=0$ and
$B(s^{*})=H+T[s^{*}]+\sigma\|s^{*}\|^{2}I_{n}$ ${-\nu\big(\,\|s^{*}\|^{2}I_{n}-s^{*}s^{*T}\big)}-\frac{2}{\sigma-\nu}\,tt^{T}-2(s^{*}t^{T}+ts^{*T})$
$\qquad-\sum_{1\leq i<j\leq n}\left(\tilde{s}^{(ij)}b^{(ij)T}+b^{(ij)}\tilde{s}^{(ij)T}+\frac{1}{\nu}b^{(ij)}b^{(ij)T}\right)\succeq 0.$ (9)
where $\tilde{s}^{(ij)}=[0,\dotsc s^{*}_{j},\dotsc,s^{*}_{i},\dotsc 0]\in\mathbb{R}^{n}$ has $s^{*}_{j}$ on $i$th entry and $s^{*}_{i}$ on $j$th entry and all other entries are zero; $b^{(ij)}$ and $t$ depend on tensor coefficients and are defined in (8). Then, $q(v):=m_{3}(s^{*}+v)-m_{3}(s^{*})$ is SoS in $v$, and consequently, $s^{*}$ is a global minimizer of $m_{3}$.
Proof.
Since $\nabla m_{3}(s^{*})=0$, from Lemma 2.1,
$2q(v)=\nabla^{2}m_{3}(s^{*})[v]^{2}+\frac{1}{3}T[v]^{3}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2\sigma\|v\|^{2}v^{T}s^{*}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\sigma}{2}\|v\|^{4}}.$ (10)
We aim to express every term in (10) with $\omega=[1,v,u,z]$. Fix a constant $\nu>0$, we write
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sigma v^{T}s^{*}\|v\|^{2}}$ $=$ $(\sigma-\nu)v^{T}{s^{*}}\|v\|^{2}+\nu\sum_{i=1}^{n}{s^{*}_{i}v_{i}^{3}}+\nu\sum_{1\leq i<j\leq n}v_{i}v_{j}(v_{i}s^{*}_{j}+v_{j}s^{*}_{i})$ (11)
$=$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(\sigma-\nu)\bigg(\sum_{i=1}^{n}u_{i}\bigg)(v^{T}{s^{*}})+\nu\sum_{i=1}^{n}{s^{*}_{i}v_{i}u_{i}}+\nu\sum_{1\leq i<j\leq n}z_{ij}(v^{T}\tilde{s}^{(ij)})}$
where $\tilde{s}^{(ij)}=[0,\dotsc s^{*}_{j},\dotsc,s^{*}_{i},\dotsc 0]\in\mathbb{R}^{n}$ has $s^{*}_{j}$ on $i$th entry and $s^{*}_{i}$ on $j$th entry and all other zeros. Also,
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma\|v\|^{4}}=\sigma\bigg(\sum_{i=1}^{n}v_{i}^{2}\bigg)^{2}=\sigma\bigg(\sum_{i=1}^{n}v_{i}^{4}+2\sum_{1\leq i<j\leq n}v_{i}^{2}v_{j}^{2}\bigg)={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma\sum_{i=1}^{n}u_{i}^{2}+2(\sigma-\nu)\sum_{1\leq i<j\leq n}u_{i}u_{j}+2\nu\sum_{1\leq i<j\leq n}z_{ij}^{2}.}$ (12)
Following from (8), (10), (11), (12), we have
$2q(v)$ $=$ $\nabla^{2}m_{3}(s^{*})[v]^{2}+\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\sigma-\nu}{2}\sum_{i=1}^{n}u_{i}^{2}+(\sigma-\nu)\sum_{1\leq i<j\leq n}u_{i}u_{j}}+2\bigg(\sum_{i=1}^{n}u_{i}\bigg)\big(t^{T}v\big)+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2(\sigma-\nu)\bigg(\sum_{i=1}^{n}u_{i}\bigg)(v^{T}{s^{*}})}}_{:=\mathcal{I}_{1}}$ (13)
$\hskip 18.49988pt+\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\nu\sum_{1\leq i<j\leq n}z_{ij}^{2}}+\sum_{1\leq i<j\leq n}2z_{ij}\big({v^{T}b^{(ij)}}\big)+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\nu\sum_{1\leq i<j\leq n}2z_{ij}(v^{T}\tilde{s}^{(ij)})}}_{:=\mathcal{I}_{2}}+\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2\nu\sum_{i=1}^{n}{s^{*}_{i}v_{i}u_{i}}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\nu}{2}\sum_{i=1}^{n}u_{i}^{2}}}_{:=\mathcal{I}_{3}}.$
Note that the expressions in blue represents the homogeneous quartic terms $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\sigma}{2}\|v\|^{4}$} from (12). The expressions in red represents $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sigma v^{T}s^{*}\|v\|^{2}$} from (11) and the rest of the cubic terms represents the tensor $\frac{1}{3}T[v]^{3}$ from (8). $\mathcal{I}_{1}$, $\mathcal{I}_{2}$, and $\mathcal{I}_{3}$ can be reorganized as
$\mathcal{I}_{1}$ $=$ $\underbrace{\frac{\sigma-\nu}{2}\bigg[\sum_{i=1}^{n}u_{i}+2(v^{T}s^{*})+\frac{2t^{T}v}{\sigma-\nu}\bigg]^{2}}_{:={\mathcal{SI}_{1}}}-\underbrace{\bigg(\frac{2}{\sigma-\nu}(t^{T}v)^{2}+4(v^{T}s^{*})(t^{T}v)\bigg)}_{\mathcal{BI}_{1}}-{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}2(\sigma-\nu)(v^{T}s^{*})^{2}}$
$\mathcal{I}_{2}$ $=$ $\underbrace{\nu\sum_{1\leq i<j\leq n}\bigg[z_{ij}+v^{T}\tilde{s}^{(ij)}+\frac{v^{T}b^{(ij)}}{\nu}\bigg]^{2}}_{:={\mathcal{SI}_{2}}}-\underbrace{\sum_{1\leq i<j\leq n}\bigg[2(v^{T}\tilde{s}^{(ij)})(v^{T}b^{(ij)})+\frac{(v^{T}b^{(ij)})^{2}}{\nu}}_{\mathcal{BI}_{2}}+{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\nu(v^{T}\tilde{s}^{(ij)})^{2}}\bigg]$
$\mathcal{I}_{3}$ $=$ $\underbrace{\frac{\nu}{2}\sum_{i=1}^{n}\big(u_{i}+2s_{i}^{*}v_{i}\big)^{2}}_{:={\mathcal{SI}_{3}}}-{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}2\nu\sum_{i=1}^{n}(s_{i}^{*}v_{i})^{2}}.$
We collect the terms that contain multiple of $v_{i}$ and $s_{i}^{*}$, i.e., the last term in $\mathcal{I}_{1}$, $\mathcal{I}_{2}$, and $\mathcal{I}_{3}$, respectively (highlighted in purple),
$2(\sigma-\nu)(v^{T}s^{*})^{2}+\nu\sum_{1\leq i<j\leq n}(v_{i}s_{j}^{*}+v_{j}s_{i}^{*})^{2}+2\nu\sum_{i=1}^{n}(s_{i}^{*}v_{i})^{2}$ (14)
$=$ $2(\sigma-\nu)(v^{T}s^{*})^{2}+\nu\sum_{1\leq i<j\leq n}(v_{i}s_{j}^{*}+v_{j}s_{i}^{*})^{2}+2\nu(v^{T}s^{*})^{2}-4\nu\sum_{1\leq i<j\leq n}s_{i}^{*}s_{j}^{*}v_{i}v_{j}$
$=$ $2\sigma(v^{T}s^{*})^{2}+\underbrace{\nu\sum_{1\leq i<j\leq n}(v_{i}s_{j}^{*}-v_{j}s_{i}^{*})^{2}}_{:={\mathcal{BI}_{3}}}$
Note that
$\mathcal{BI}_{3}$ $=$ $\nu\sum_{1\leq i<j\leq n}v_{i}^{2}{s_{j}^{*}}^{2}+v_{j}^{2}{s_{i}^{*}}^{2}-2s_{i}^{*}s_{j}^{*}v_{i}v_{j}=\nu\bigg(\sum_{i=1}^{n}v_{i}^{2}\bigg)\|s^{*}\|^{2}-\nu\bigg(\sum_{i=1}^{n}s_{i}^{*}v_{i}\bigg)^{2}=\nu v^{T}\bigg(\|s^{*}\|^{2}I_{n}-s^{*}{s^{*}}^{T}\bigg)v.$
$\mathcal{BI}_{1}$ $=$ $\frac{2}{\sigma-\nu}(t^{T}v)^{2}+4(v^{T}s^{*})(t^{T}v)=v^{T}\bigg[\frac{2\,tt^{T}}{\sigma-\nu}+2\big(s^{*}t^{T}+t{s^{*}}^{T}\big)\bigg]v.$
$\mathcal{BI}_{2}$ $=$ $\sum_{1\leq i<j\leq n}\bigg[2(v^{T}\tilde{s}^{(ij)})(v^{T}b^{(ij)})+\frac{(v^{T}b^{(ij)})^{2}}{\nu}\bigg]=v^{T}\bigg[\big(\tilde{s}^{(ij)}b^{(ij)}{}^{T}+b^{(ij)}\tilde{s}^{(ij)}{}^{T}\big)+\frac{b^{(ij)}{b^{(ij)}}^{T}}{\nu}\bigg]v.$
Using $\mathcal{I}_{1}$, $\mathcal{I}_{2}$, $\mathcal{I}_{3}$ , (14) and $\mathcal{BI}_{1}$, $\mathcal{BI}_{2}$, and $\mathcal{BI}_{3}$ ,
$2q(v)$ $=$ $\bigg(\nabla^{2}m_{3}(s^{*})[v]^{2}-2\sigma(v^{T}s^{*})^{2}-{\mathcal{BI}_{1}}-{\mathcal{BI}_{2}}-{\mathcal{BI}_{3}}\bigg)+{\mathcal{SI}_{1}}+{\mathcal{SI}_{2}}+{\mathcal{SI}_{3}}$ (15)
$=$ $B(s^{*})[v]^{2}+{\mathcal{SI}_{1}}+{\mathcal{SI}_{2}}+{\mathcal{SI}_{3}}$
where $\mathcal{SI}_{1}$} to $\mathcal{SI}_{3}$} are the nonnegative square polynomials in $v$ defined in $\mathcal{I}_{1}$ to $\mathcal{I}_{3}$. Comparing our derivations with the expression of $B(s^{*})$ in (2.1), we see that the first three terms of $B(s^{*})$ come from $\nabla^{2}m_{3}(s^{*})-2\sigma s^{*}{s^{*}}^{T}$, while the remaining terms in $B(s^{*})$ are from $\mathcal{BI}_{1}$} to $\mathcal{BI}_{3}$}. The matrix $B(s^{*})$ depends only on $s^{*}$ and the coefficients of the model $m_{3}$, namely $g$, $H$, $T$, and $\sigma$. Since $B(s^{*})$ is symmetric and positive semidefinite, the quadratic form $B(s^{*})[v]^{2}$ is a sum of squares. Together with the remaining sum-of-squares terms ${\mathcal{SI}_{1}}+{\mathcal{SI}_{2}}+{\mathcal{SI}_{3}}$, this implies that $q(v)$ is itself a sum-of-squares polynomial. Since $q(v):=m_{3}(s^{*}+v)-m_{3}(s^{*})\geq 0$ for all $v$, we deduce that $s^{*}$ is a global minimizer. ∎
Remark 2.2.
The SoS certificate condition established in Theorem 2.1 is closely related to global optimality conditions. For example, when $T=0$, by [17, 22], a point $s^{*}$ is a global minimizer of $m_{3}$ if and only if
$\nabla m_{3}(s^{*})=0,\qquad B(s^{*})=H+\sigma\|s^{*}\|^{2}I_{n}\succeq 0,$ (16)
which coincides with (2.1). Yet in the case of $T\neq 0$, the condition in (2.1) does not, in general, coincide with the global optimality conditions for $m_{3}$ derived in [76, Sec. 3].)
Though Theorem 2.1 implies that $s^{*}$ is a local minimizer, we ‘double check’ our results by showing this fact in detail.
Lemma 2.3.
(SoS certificate implies local optimality conditions) The conditions in Theorem 2.1 imply that both first- and second-order local optimality conditions of $m_{3}$ are satisfied at $s^{*}$; namely, $\nabla m_{3}(s^{*})=0$ and $\nabla^{2}m_{3}(s^{*})\succeq 0$, respectively.
Proof.
The condition $\nabla m_{3}(s^{*})=0$ in Theorem 2.1 already gives the first-local optimality conditions of $m_{3}$. It remains to show that $B(s^{*})\succeq 0$ implies $\nabla^{2}m_{3}(s^{*})\succeq 0$. Note that $\nabla^{2}m_{3}(s^{*})=H+T[s^{*}]+\sigma\|s^{*}\|^{2}I_{n}+2\sigma\,s^{*}{s^{*}}^{T}$. We rearrange (2.1) into
$\nabla^{2}m_{3}(s^{*})$ $\succeq$ $2\sigma\,s^{*}{s^{*}}^{T}+\nu\big(\,\|s^{*}\|^{2}I_{n}-s^{*}s^{*T}\big)+$ (17)
$\qquad\underbrace{\frac{2}{\sigma-\nu}\,tt^{T}+2(s^{*}t^{T}+ts^{*T})}_{:=\mathcal{BF}_{1}}+\underbrace{\sum_{1\leq i<j\leq n}\left(\tilde{s}^{(ij)}b^{(ij)T}+b^{(ij)}\tilde{s}^{(ij)T}+\frac{1}{\nu}b^{(ij)}b^{(ij)T}\right)}_{:=\mathcal{BF}_{2}}.$
We have
$\mathcal{BF}_{1}=2\bigg(\sqrt{(\sigma-\nu)^{-1}}t+\sqrt{\sigma-\nu}s^{*}\bigg)\bigg(\sqrt{(\sigma-\nu)^{-1}}t+\sqrt{\sigma-\nu}s^{*}\bigg)^{T}-2(\sigma-\nu)s^{*}{s^{*}}^{T}$ (18)
and
$\mathcal{BF}_{2}=\sum_{1\leq i<j\leq n}\bigg(\sqrt{{\nu}}\tilde{s}^{(ij)}+\sqrt{{\nu}^{-1}}b^{(ij)}\bigg)\bigg(\sqrt{{\nu}}\tilde{s}^{(ij)}+\sqrt{{\nu}^{-1}}b^{(ij)}\bigg)^{T}-\nu\sum_{1\leq i<j\leq n}\tilde{s}^{(ij)}\tilde{s}^{(ij)T}.$ (19)
Since $\tilde{s}^{(ij)}=[0,\dotsc s^{*}_{j},\dotsc,s^{*}_{i},\dotsc 0]\in\mathbb{R}^{n}$, we can write $\tilde{s}^{(ij)}:=s^{*}_{j}e_{i}+s^{*}_{i}e_{j},$ where $\{e_{k}\}_{k=1}^{n}$ denotes the standard basis of $\mathbb{R}^{n}$. Then $\tilde{s}^{(ij)}\tilde{s}^{(ij)T}={s^{*}_{j}}^{2}\,e_{i}e_{i}^{T}+{s^{*}_{i}}^{2}\,e_{j}e_{j}^{T}+s^{*}_{i}s^{*}_{j}\big(e_{i}e_{j}^{T}+e_{j}e_{i}^{T}\big).$ Summing over all $1\leq i<j\leq n$, the resulting matrix has diagonal entries $\sum_{\ell\neq k}s_{\ell}^{2}=\|s^{*}\|^{2}-s_{k}^{2},$ and off-diagonal entries $M_{pq}=s^{*}_{p}s^{*}_{q}\quad(p\neq q).$ The sum admits the compact representation
$\sum_{1\leq i<j\leq n}\tilde{s}^{(ij)}\tilde{s}^{(ij)T}=\|s^{*}\|^{2}I+s^{*}{s^{*}}^{T}-2\,\operatorname{diag}\{{s^{*}_{1}}^{2},\dots,{s^{*}_{n}}^{2}\}.$ (20)
Substituting (18), (19), (20) in (17), we obtain that
$\nabla^{2}m_{3}(s^{*})$ $\succeq$ $\underbrace{2\sigma\,s^{*}{s^{*}}^{T}+\nu\big(\,\|s^{*}\|^{2}I_{n}-s^{*}s^{*T}\big)-2(\sigma-\nu)s^{*}{s^{*}}^{T}-\nu\bigg(\|s^{*}\|^{2}I+s^{*}{s^{*}}^{T}-2\,\operatorname{diag}\{{s^{*}_{1}}^{2},\dots,{s^{*}_{n}}^{2}\}\bigg)}_{\mathcal{BF}}$
$\qquad 2\bigg(\sqrt{(\sigma-\nu)^{-1}}t+\sqrt{\sigma-\nu}s^{*}\bigg)\bigg(\sqrt{(\sigma-\nu)^{-1}}t+\sqrt{\sigma-\nu}s^{*}\bigg)^{T}+$
$\qquad\sum_{1\leq i<j\leq n}\bigg(\sqrt{{\nu}}\tilde{s}^{(ij)}+\sqrt{{\nu}^{-1}}b^{(ij)}\bigg)\bigg(\sqrt{{\nu}}\tilde{s}^{(ij)}+\sqrt{{\nu}^{-1}}b^{(ij)}\bigg)^{T}.$
Rearranging the first term gives $\mathcal{BF}=2\nu\operatorname{diag}\!\big({s_{1}^{*}}^{2},\dots,{s_{n}^{*}}^{2}\big),$ which is positive semidefinite. The next two terms are also positive semidefinite. Hence, $\nabla^{2}m_{3}(s^{*})\succeq 0$, $B(s^{*})\succeq 0$ implies the second-order local optimality conditions of $m_{3}$. ∎
The next corollary gives a more standard SoS formulation of our result, showing that we are able to identify a $\cal Q$} matrix as needed in Theorem 1.4.
Corollary 2.1.
($\cal Q$} matrix formulation) Let $\mathbf{1}^{T}=[1,\dotsc,1]\in\mathbb{R}^{n}$. We construct a symmetric matrix of dimension $\tbinom{n}{2}\times\tbinom{n}{2}$, such that
${\cal Q}=\begin{bmatrix}0&0&0&0\\[3.0pt] 0&\nabla^{2}m_{3}(s^{*})&D(s^{*})&C(s^{*})\\[3.0pt] 0&D(s^{*})^{T}&U&0\\[3.0pt] 0&C(s^{*})^{T}&0&\nu I_{N}\end{bmatrix}$ (21)
  • The $u$-$u$ block has size $n\times n$ and $U=\Big(\frac{\sigma}{2}-\frac{\nu}{2}\Big)\mathbf{1}\mathbf{1}^{T}+\frac{\nu}{2}I_{n}$ which has $\frac{\sigma}{2}$ on the diagonal and $\frac{\sigma-\nu}{2}$ off-diagonal.
  • The $z$-$z$ block has size $N\times N$ is an identity matrix $\nu I_{N}$.
  • The $v$-$u$ block has size $n\times n$, $D(s^{*})=\big[(\sigma-\nu)s^{*}+t\big]\mathbf{1}^{T}+\nu\operatorname{diag}(s^{*})$.
  • The $v$-$z$ block has size $n\times N$ with the $ij$-th column as $\nu\tilde{s}^{(ij)}+b^{(ij)}$.
If the conditions in Theorem 2.1 are satisfied, then ${\cal Q}\succeq 0$.
Proof.
Using $\omega:=[1,v,u,z]^{T}$ as defined in (3) we analyze $\omega^{T}{\cal Q}\omega$ term by term. Using $\mathbf{1}^{T}u=\sum_{i=1}^{n}u_{i}$, the $u$-$u$ block yields
$u^{T}Uu=\frac{\sigma}{2}\sum_{i=1}^{n}u_{i}^{2}+(\sigma-\nu)\sum_{1\leq i<j\leq n}u_{i}u_{j}.$
The $z$-$z$ block gives $\nu\sum_{1\leq i<j\leq n}z_{ij}^{2}$. The $v$-$u$ block gives
$u^{T}D(s^{*})v=2\bigg(\sum_{i=1}^{n}u_{i}\bigg)\big(t^{T}v\big)+2(\sigma-\nu)\bigg(\sum_{i=1}^{n}u_{i}\bigg)(v^{T}{s^{*}})+2\nu\sum_{i=1}^{n}{s^{*}_{i}v_{i}u_{i}}.$
The $v$-$z$ block gives
$u^{T}C(s^{*})z=\nu\sum_{1\leq i<j\leq n}z_{ij}(v^{T}\tilde{s}^{(ij)})+\sum_{1\leq i<j\leq n}z_{ij}\big({v^{T}b^{(ij)}}\big).$
By adding the terms from these blocks and comparing $\omega^{T}{\cal Q}\omega$ with (13), we obtain $\omega^{T}{\cal Q}\omega=2q(v).$ If the conditions of Theorem 2.1 are satisfied, then $q(v)$ is a sum of squares. By Theorem 1.4, we conclude that ${\cal Q}\succeq 0$. ∎
Example 2.1.
We write the full matrix for the case of $n=2$:
$2q(v)=\begin{bmatrix}1\\ v_{1}\\ v_{2}\\ u_{1}\\ u_{2}\\ z_{12}\end{bmatrix}\begin{bmatrix}0&0&0&0&0&0\\ 0&\nabla^{2}m_{3}(s^{*})&&\sigma s_{1}+t_{111}&(\sigma-\nu)s_{1}+t_{111}&\nu s_{2}+t^{*}_{112}\\ 0&&&(\sigma-\nu)s_{2}+t_{222}&\sigma s_{2}+t_{222}&\nu s_{1}+t^{*}_{122}\\ 0&\ast&\ast&\frac{\sigma}{2}&\frac{\sigma}{2}-\frac{\nu}{2}&0\\ 0&\ast&\ast&\frac{\sigma}{2}-\frac{\nu}{2}&\frac{\sigma}{2}&0\\ 0&\ast&\ast&0&0&\nu\end{bmatrix}\begin{bmatrix}1\\ v_{1}\\ v_{2}\\ u_{1}\\ u_{2}\\ z_{12}\end{bmatrix}$.
The symbol “*” denotes symmetric entries in the corresponding positions. By examining the quadratic form ${\cal Q}[\tilde{\omega}]^{2}$, with $\tilde{\omega}\in\mathbb{R}^{\frac{1}{2}(n+2)(n+1)}$, we find that ${\cal Q}[\tilde{\omega}]^{2}\geq 0$ holds whenever $\nabla m_{3}(s^{*})=0$ and $B(s^{*})\succeq 0$. It follows that ${\cal Q}\succeq 0$, and consequently the corresponding polynomial $q(v)$ is SoS. Lastly, we note that the expression of $\cal Q$} is not unique. Let $\nu$, $\delta$, and $\beta$ be arbitrary constants representing the degrees of freedom in choosing the matrix $\cal Q$}.
${\cal Q}=\begin{bmatrix}0&0&0&0&0&0\\ 0&\nabla^{2}m_{3}(s^{*})&&\sigma s_{1}+t_{111}&(\sigma-\nu)s_{1}+t_{122}+\beta&\nu s_{2}+t_{112}+\delta\\ 0&&&(\sigma-\nu)s_{2}-\delta&\sigma s_{2}+t_{222}&\nu s_{1}-\beta\\ 0&\ast&\ast&\frac{\sigma}{2}&\frac{\sigma}{2}-\frac{\nu}{2}&0\\ 0&\ast&\ast&\frac{\sigma}{2}-\frac{\nu}{2}&\frac{\sigma}{2}&0\\ 0&\ast&\ast&0&0&\nu\end{bmatrix}$.
2.2 Bounds on the length of $s^{*}$
To effectively certify the positive semidefiniteness of $B(s^{*})$, we require upper and lower bounds on $s^{*}$ expressed in terms of the coefficients of $m_{3}$. Equation 29 gives these bounds. Note that we define the vector norm, the matrix norm and tensor norm as such
$\|H\|:$ $=$ $\max_{s\in\mathbb{R}^{n},\|s\|=1}|H[s]^{2}|=\max_{i}\{|\lambda_{i}|\},\qquad\lambda_{*}:=\max\{-\lambda_{\min}[H],0\}$ (22)
$\|T\|:$ $=$ $\max_{s\in\mathbb{R}^{n},\|s\|=1}|T[s]^{3}|,\qquad\|T\|_{F}:=\sum_{i,j,k}T_{ijk}^{2},\qquad\|g\|=(\sum_{i=1}^{n}g_{i}^{2})^{1/2}.$ (23)
Clearly, $\|T\|_{F}\geq\|T\|$ and more details are provided in Section A.1.
Theorem 2.2.
(Bound for $s^{*}$) Assume that $\|g\|\neq 0$, let $s^{*}$ satisfies $\nabla m_{3}(s^{*})=0$ and $m_{3}(s^{*})\leq m_{3}(0)$, and $\sigma$ satisfies
$\sigma\geq\max\bigg\{\frac{9\|H\|^{3}}{\|g\|^{2}},\bigg(\frac{3\|T\|^{3}}{8\|g\|}\bigg)^{1/2}\bigg\},$ (24)
then,
$\bigg(\frac{\|g\|}{3\sigma}\bigg)^{1/3}\leq\|s^{*}\|\leq 2\bigg(\frac{\|g\|}{\sigma}\bigg)^{1/3}.$ (25)
Proof.
To prove the upper bound for $\|s^{*}\|$, we rewrite $m_{3}$ at $s^{*}$ as
$m_{3}(s^{*})-m_{3}(0)$ $=$ $\bigg(g^{T}s^{*}+\frac{\sigma}{8}\|s^{*}\|^{4}\bigg)+\bigg(\frac{1}{2}H[s^{*}]^{2}+\frac{\sigma}{16}\|s^{*}\|^{4}\bigg)+\bigg(\frac{1}{6}T[s^{*}]^{3}+\frac{\sigma}{16}\|s^{*}\|^{4}\bigg)$
$\geq$ $\bigg(-\|g\|+\frac{\sigma}{8}\|s^{*}\|^{3}\bigg)\|s^{*}\|+\bigg(-\frac{\lambda_{*}}{2}+\frac{\sigma}{16}\|s^{*}\|^{2}\bigg)\|s^{*}\|^{2}+\bigg(-\frac{\|T\|}{6}+\frac{\sigma}{16}\|s^{*}\|\bigg)\|s^{*}\|^{3}$
where the last inequality comes from $g^{T}s\leq\|g\|\|s\|$, $H[s^{*}]^{2}\geq-\lambda_{*}\|s^{*}\|$ and $T[s^{*}]^{3}\geq-\|T\|\|s^{*}\|^{3}$. Assume that (24) holds and assume for contradiction the global minimum of $m_{3}(s)$ happens at $\|s^{*}\|>r_{c}:=\big(\frac{8\|g\|}{\sigma}\big)^{1/3}$. Then,
$m_{3}(s^{*})-m_{3}(0)>\underbrace{\bigg[-\|g\|+\frac{\sigma}{8}r_{c}^{3}\bigg]}_{=0}\|s^{*}\|+\underbrace{\bigg[-\frac{\lambda_{*}}{2}+\frac{\sigma}{16}r_{c}^{2}\bigg]}_{\geq 0}\|s^{*}\|^{2}+\underbrace{\bigg[-\frac{\|T\|}{6}+\frac{\sigma}{16}r_{c}\bigg]}_{\geq 0}\|s^{*}\|^{3}\geq 0.$ (26)
However, we have $m_{3}(s^{*})\leq m_{3}(0)$, we arrives at a contradiction. Note that in (26), the first term is zero using $r_{c}=\big(\frac{8\|g\|}{\sigma}\big)^{1/3}$. The second and third term is positive using $\sigma\geq\frac{9\|H\|^{3}}{\|g\|^{2}}$ and $\sigma\geq\big(\frac{3\|T\|^{3}}{8\|g\|}\big)^{1/2}$ from condition (24),
[…]
To prove the lower bound for $\|s^{*}\|$,
$\big\|g\big\|$ $=$ $\big\|\nabla m_{3}(0)\big\|=\big\|\nabla m_{3}(s^{*})-\nabla m_{3}(0)\big\|\leq\bigg\|H[s^{*}]+\frac{1}{2}T[s^{*}]^{2}+\sigma\big\|s^{*}\big\|^{2}s^{*}\bigg\|$ (27)
$\leq$ $\|H\|\|s^{*}\|+\frac{1}{2}\|T\|\|s^{*}\|^{2}+\sigma\|s^{*}\|^{3}.$
Therefore, we have
$\frac{\|g\|}{3}\leq\|H\|\|s^{*}\big\|,\quad\text{or}\quad\frac{\|g\|}{3}\leq\frac{1}{2}\|T\|\|s^{*}\|^{2},\quad\text{or}\quad\frac{\|g\|}{3}\leq\sigma\|s^{*}\|^{3}.$
This gives
$\|s^{*}\|\geq\min\bigg\{\frac{\|g\|}{3\|H\|},\bigg(\frac{2\|g\|}{3\|T\|}\bigg)^{1/2},\bigg(\frac{\|g\|}{3\sigma}\bigg)^{1/3}\bigg\}.$ (28)
From (24), we have
$\sigma\geq\frac{9\|H\|^{3}}{\|g\|^{2}},\qquad$ $\Rightarrow$ $\qquad\bigg(\frac{\|g\|}{3\sigma}\bigg)^{1/3}\leq\frac{\|g\|}{3\|H\|}$
$\sigma\geq\bigg(\frac{3\|T\|^{3}}{8\|g\|}\bigg)^{1/2}\qquad$ $\Rightarrow$ $\qquad\bigg(\frac{\|g\|}{3\sigma}\bigg)^{1/3}\leq\bigg(\frac{2\|g\|}{3\|T\|}\bigg)^{1/2}.$
Therefore, under (24), (28) simplifies to $\|s^{*}\|\geq\big(\frac{\|g\|}{3\sigma}\big)^{1/3}.$
Remark 2.3.
(Global Minimum Attained) For any $\sigma>0$, the model $m_{3}(s)$ is bounded below and coercive and $\|s\|\to\infty$ gives $m_{3}(s)\to+\infty.$ The argument used to derive an upper bound on $\|s^{*}\|$ implies that $m_{3}(s)\leq m_{3}(0)$ can only occur when $\|s\|\leq r_{c}$ for some positive $r_{c}$. Consequently, the constrained minimization problem $s_{r}^{*}:=\arg\min_{\|s\|\leq r_{c}}m_{3}(s)$ admits a global minimizer (not just a finite global minimum). Since any global minimizer $s^{*}$ of $m_{3}$ must satisfy $m_{3}(s^{*})\leq m_{3}(0)$, we conclude that $\|s^{*}\|\leq r_{c}$, and hence a global minimizer of $m_{3}$ is attained for any $\sigma>0$.
2.3 Behaviour of $B(s^{*})$ when $\sigma$ large
We are now ready to prove the main theoretical result of the paper. In Theorem 2.3, we show that when $\sigma$ is sufficiently large, the matrix $B(s^{*})$ becomes positive semidefinite, and consequently $m_{3}(s^{*}+v)-m_{3}(s^{*})$ is a SoS polynomial. Moreover, we provide an explicit bound on $\sigma$, expressed solely in terms of the coefficients of $m_{3}$, that guarantees the SoS representation of $m_{3}(s^{*}+v)-m_{3}(s^{*})$. In Section 2.3.1 and Section 2.3.2, we provide a case-by-case analysis showing how the bound simplifies for locally convex and nonconvex $m_{3}$.
Theorem 2.3.
($B(s^{*})$ is PSD for large $\sigma$) Assume that $\|g\|\neq 0$, let $s^{*}$ satisfies $\nabla m_{3}(s^{*})=0$ and $m_{3}(s^{*})\leq m_{3}(0)$, and $\sigma$ satisfies
${{\sigma\geq\max\bigg\{\frac{9\|H\|^{3}}{\|g\|^{2}},\bigg(\frac{3\|T\|^{3}}{8\|g\|}\bigg)^{1/2},\bigg(\frac{3\cdot 32^{3}\|T\|^{3}}{\|g\|}\bigg)^{1/2},\frac{3^{2}\cdot 6^{3}\lambda_{*}^{3}}{\|g\|^{2}}\bigg\}=\max\bigg\{\frac{9\cdot 6^{3}\|H\|^{3}}{\|g\|^{2}},\bigg(\frac{3\cdot 32^{3}\|T\|^{3}}{\|g\|}\bigg)^{1/2}\bigg\}}}$ (29)
where the norms and $\lambda_{*}$ are defined in (22) – (23). Then $B(s^{*})$ defined in (2.1) with $\nu=\frac{\sigma}{2}$ is positive semi-definite. Moreover, then, $m_{3}(s^{*}+v)-m_{3}(s^{*})$ is SoS and consequently, $s^{*}$ is a global minimizer.
Proof.
When $\nu=\frac{\sigma}{2}$ in (2.1), we have
$B(s^{*})=H+T[s^{*}]+\frac{\sigma}{2}\|s^{*}\|^{2}I_{n}+\frac{\sigma}{2}s^{*}{s^{*}}^{T}-\frac{4tt^{T}}{\sigma}-2\big(s^{*}t^{T}+t{s^{*}}^{T}\big)-\sum_{1\leq i<j\leq n}\Big[\tilde{s}^{(ij)}b^{(ij)T}+b^{(ij)}\tilde{s}^{(ij)T}+\frac{2b^{(ij)}b^{(ij)T}}{\sigma}\Big]$
We give some lower bounds for the eigenvalues of the terms in $B(s^{*})$, using the inequalities
${-2\|t\|\cdot\|s^{*}\|I_{n}\preceq}\,\,s^{*}t^{T}+t{s^{*}}^{T}$ $\preceq$ $2\|t\|\cdot\|s^{*}\|I_{n},\qquad 0\preceq tt^{T}\preceq\|t\|^{2}I_{n},$ (30)
$0\preceq\sum_{1\leq i<j\leq n}b^{(ij)}b^{(ij)T}$ $\preceq$ $\sum_{1\leq i<j\leq n}\|b^{(ij)}\|^{2}I_{n}\preceq\|T\|_{F}^{2}I_{n},\quad 0\preceq s^{*}{s^{*}}^{T},$
${-2\|T\|_{F}\|s^{*}\|I_{n}\preceq}\sum_{1\leq i<j\leq n}\Big[\tilde{s}^{(ij)}b^{(ij)T}+b^{(ij)}\tilde{s}^{(ij)T}\Big]$ $\preceq$ $2\sum_{1\leq i<j\leq n}\|\tilde{s}^{(ij)}\|\cdot\|b^{(ij)T}\|I_{n}\preceq 2\|T\|_{F}\|s^{*}\|I_{n}.$
Using the upper bounds in (30), and definition of norms in (22) – (23), we have
$\lambda_{\min}\big[B(s^{*})\big]$ $\geq$ $-\lambda_{*}-\|T\|\|s^{*}\|+\frac{\sigma}{2}\|s^{*}\|^{2}-\frac{4}{\sigma}\|t\|^{2}-4\|t\|\|s^{*}\|-2\|T\|_{F}\|s^{*}\|-\frac{2}{\sigma}\|T\|_{F}^{2},$ (31)
$\geq$ $\frac{\sigma}{2}\|s^{*}\|^{2}-7\|T\|_{F}\|s^{*}\|-\frac{6}{\sigma}\|T\|^{2}_{F}-\lambda_{*},$ (32)
where in the second inequality, we used $\|T\|_{F}\geq\|T\|$ and $\|T\|_{F}\geq\|t\|$, and re-arrangements. Thus, $\lambda_{\min}\big[B(s^{*})\big]\geq 0$ if the right-hand side of (32) is non-negative, which in turn is equivalent to $\|s_{*}\|$ being greater or equal to the positive root of the right-hand side of (32) (as a function of $\|s_{*}\|$), namely, $\|s^{*}\|\geq\frac{7}{\sigma}\|T\|_{F}+\frac{1}{\sigma}\sqrt{2\sigma\lambda_{*}+61\|T\|^{2}_{F}}$. In particular, the latter is achieved if $\|s^{*}\|\geq\max\left\{\frac{32}{\sigma}\|T\|_{F},6\left(\frac{\lambda_{*}}{\sigma}\right)^{1/2}\right\}$. We now recall the lower bound in (25), namely, $\|s^{*}\|\geq\big(\frac{\|g\|}{3\sigma}\big)^{1/3}$, which we easily check that it satisfies $\big(\frac{\|g\|}{3\sigma}\big)^{1/3}\geq\max\left\{\frac{32}{\sigma}\|T\|_{F},6\left(\frac{\lambda_{*}}{\sigma}\right)^{1/2}\right\}$ when $\sigma$ satisfies (29).
The bound in (29) contains $\|g\|$ in the denominator. When the quartically regularized polynomial is used as a subproblem within a higher-order method, the stopping condition is typically chosen as $\|g\|\geq\epsilon_{g}$, meaning that the algorithm terminates once an approximate critical point of the subproblem has been identified. As shown in [19], even an approximate critical point is sufficient to guarantee convergence and complexity of the outer algorithm. Therefore, the bound in (29) is well defined in the context of high-order optimization methods. For completeness, we also discuss the behaviour in the degenerate case $\|g\|=0$ in Remark 2.4.
Remark 2.4.
We give a discussion on what would happen if $\|g\|=0$.
  • If $H\succeq\delta I_{n}\succ 0$, then $m_{3}(s)-m_{3}(s^{*})$ is both SoS and SoS-convex for any $\sigma>\tfrac{8\|T\|^{2}}{\delta}$; see Section 2.3.1.
  • If $T=0$, or if $T$ has special structure $T[s]^{3}=t^{T}s\,\|s\|^{2}$, then $m_{3}(s)-m_{3}(s^{*})$ is SoS for every $\sigma>0$; see Section 3.2.
  • If $H$ is not PSD and $T$ is general, determining whether $m_{3}(s)-m_{3}(s^{*})$ becomes SoS for sufficiently large $\sigma>0$ remains an open question. We leave this analysis to future work.
Beyond the behavior of $B(s^{*})$, we also remark that if $s^{*}$ is any second-order local minimizer of the model $m_{3}(s)$, then $\|s^{*}(\sigma)\|$ is a non-increasing function of $\sigma$. Further details are provided in Section A.2.
2.3.1 Locally Strongly Convex $m_{3}$: Convexification, SoS convex and Unique global minimizer
If $m_{3}$ is locally strongly convex at $s=0$, i.e., $H=\nabla^{2}m_{3}(0)\succeq\delta I_{n}\succ 0$, then for sufficiently large $\sigma$, the model $m_{3}$ remains strongly convex and is, in fact, SoS convex for all $s\in\mathbb{R}^{n}$ [1]. A detailed discussion on how convexification relates to SoS and SoS-convexity can be found in [76, Sec. 2.3.2]. Here we highlight the key arguments.
If $m_{3}$ is strongly convex at $s=0$, 994 then  [76, Lemma 2.3] shows that choosing $\sigma$ such that
$\sigma>\frac{\|T\|^{2}}{4\delta}$ (33)
guarantees $m_{3}$ is strongly convex for all $s\in\mathbb{R}^{n}$. A sufficient condition for a unique global minimizer is the global strong convexity of $m_{3}$, i.e., $\nabla^{2}m_{3}(s)\succeq\delta I_{n}$ for all $s$ and some $\delta>0$. Then, (33) ensures $m_{3}$ has a unique global minimizer. Moreover, if we increase $\sigma$ further so that
$\sigma>\frac{8\|T\|^{2}}{\delta},$ (34)
then by [1], $m_{3}$ is SoS-convex. Then, it follows from Lemma 1.1 that $q(s)=m_{3}(s)-m_{3}(s^{*})$ is SoS.
2.3.2 Locally Nonconvex $m_{3}$: Global Optimality and Size of $\sigma$
The size of $\sigma$ is closely related to the idea of convexification of $m_{3}$ with large $\sigma$. If $m_{3}$ is locally nonconvex at $s=0$, then no matter how large $\sigma>0$ is chosen, it is not possible to make $m_{3}$ globally convex for all $s\in\mathbb{R}^{n}$. However, by selecting $\sigma$ sufficiently large, $m_{3}$ becomes convex everywhere except within a small neighborhood of the origin. In this section, we show that for sufficiently large $\sigma$, the stationary points (may be non-unique) lie outside the small locally nonconvex region and are global minimizers of the nonconvex model $m_{3}$. Moreover, in this regime, the global necessary and sufficient optimality conditions established in [76] coincide (more details are discussed [76, Sec 2.3.1] and Section A.3). $m_{3}(s)$ is SoS exact with zero levels of relaxation. If the global minimizer is unique, the minimum can then be found in polynomial time.
Lemma 2.4.
(Lemma 2.4 from [76]) If $\lambda_{\min}[H]\leq 0$, let $s_{0}>0$ be any scalar, and suppose that $\sigma$ satisfies
$\sigma>2\max\bigg\{\lambda_{*}s_{0}^{-2},\quad\|T\|s_{0}^{-1}\bigg\},$ (35)
then $m_{3}(s)$ is convex for all $\|s\|\geq s_{0}$.
Corollary 2.2.
Assume that $m_{3}$ is locally nonconvex at $s=0$, that is, $H=\nabla^{2}m_{3}(0)$ is not positive definite. Let $s^{*}\in\mathbb{R}^{n}$ satisfies $\nabla m_{3}(s^{*})=0$. If $\sigma$ large enough such that (63) holds, then $m_{3}(s)$ is convex for all $\|s\|\geq\sqrt{\frac{2}{3}}\,\|s^{*}\|$.
Proof.
Suppose that (63) holds. Then, by Result 1 in Theorem A.2, we have
$\sigma\geq 2\max\bigg\{\lambda_{*}\Big\|\sqrt{\frac{2}{3}}\,s^{*}\Big\|^{-2},\|T\|\,\Big\|\frac{2}{3}s^{*}\Big\|^{-1}\bigg\}>2\max\bigg\{\lambda_{*}\|s_{0}\|^{-2},\|T\|\,\|s_{0}\|^{-1}\bigg\},$
where $s_{0}=\sqrt{\frac{2}{3}}\,s^{*}$. Therefore, by Lemma 2.4, $m_{3}(s)$ is convex for all $\|s\|\geq\|s_{0}\|=\sqrt{\frac{2}{3}}\,\|s^{*}\|$. ∎
Remark 2.5 (Region of convexification).
If $\sigma$ is sufficiently large so that (29) holds, then any stationary point $s_{*}$ of $m_{3}$ with $m_{3}(s_{*})\leq m_{3}(0)$ is a global minimizer (Theorem 2.3 and Theorem 2.1). In this regime, $m_{3}$ is nonconvex only in a small neighborhood of the origin and is convex outside the region $\{\,s\in\mathbb{R}^{n}:\|s\|\leq\|s_{0}\|=\sqrt{\tfrac{2}{3}}\,\|s^{*}\|\,\}.$ Moreover, any global minimizer lies outside this nonconvex region, and no stationary point with $m_{3}(s_{*})\leq m_{3}(0)$ exists within it. This follows because, when (63) holds, every stationary point of $m_{3}$ with $m_{3}(s_{*})\leq m_{3}(0)$ is a global minimizer, and the global minimizer necessarily lies in the convex region, implying that no stationary point with $m_{3}(s_{*})\leq m_{3}(0)$ can exist within it.
Remark 2.6.
(Condition for uniqueness of the global minimizer) As shown in this section, there are no stationary points with $m_{3}(s_{*})\leq m_{3}(0)$ inside the inner nonconvex region. To establish the uniqueness of the global minimizer, one must additionally guarantee that the stationary point outside the nonconvex region is unique. In that case, this unique stationary point is also the unique global minimizer. Without such additional conditions, we can ensure existence of a global minimizer, but not necessarily uniqueness.
Remark 2.7.
(Comparing different $\sigma$ bound) [76, Sec. 2.3.1] (also summarized in this paper’s notation in Theorem A.2) provides a bound under which the necessary and sufficient conditions coincide. The bounds in Theorem A.2 and Theorem 2.3 are comparable in magnitude; both depend only on the coefficients of $m_{3}$, and both are of order [^3] In iterative algorithms for minimizing $f$, We solve a sequence of cubic subproblems $m_{3}$, with a stopping criterion $\|g\|\geq\epsilon_{g}$. Therefore, in both bounds we have $\sigma=O(\epsilon_{g}^{-2})$. $\mathcal{O}(\|g\|^{-2})$. In the locally nonconvex regime, a stationary point that achieves a decrease in the model value is the global minimizer, and the necessary and sufficient conditions coincide. Since $m_{3}(s)-m_{3}(s^{*})$ is a SoS polynomial, if the minimizer is unique, the global minimization of $m_{3}$ can be reformulated as a semidefinite program (SDP), from which the global minimum can be recovered in polynomial time.
3 Classes of SoS Quartically Regularized Polynomials
In this section, we present several special cases and variants of the quartically regularized cubic model. In some of these cases, it is possible to remove the lower bound on $\sigma$ in (29), and still ensure that $m_{3}(s)-m_{3}(s^{*})$ is SoS for all $\sigma>0$ where $s^{*}$ denotes the global minimizer. Namely,
  • Univariate quartically regularized cubic model ($n=1$), see Section 3.1.
  • Quadratic quartically regularized model (QQR), where the cubic tensor is zero ($T=0$), see Theorem 3.3.
  • A special tensor model where $T[s]^{3}=(t^{T}s)\,\|s\|^{2}$ with $t\in\mathbb{R}^{n}$, see Section 3.2.
We also investigate the Schnabel model introduced in [66] for unconstrained optimization (see Section 3.3). The model is based on a fourth-order regularization term involving the separable quartic norm $\sum_{i=1}^{n}s_{i}^{4}$. For this separable quartic norm, we find that there exist quartic polynomials that cannot be expressed as SoS, even in the limit as $\sigma\to\infty$. This observation is consistent with the example given in [2, Thm 3.3]. A summary of all cases is given in Table 1.
Table 1: Summary of SoS conditions for $q(v)=m_{3}(s^{*}+v)-m_{3}(s^{*})$ and their relation to global optimality.
Polynomial Type Condition for $q(v)$ to be SoS Is $m_{3}(s)-m_{3}(s^{*})$ SoS?
General $m_{3}$ (Section 2.3) $\nabla m_{3}(s^{*})=0$ and $B(s^{*})\succeq 0$ (2.1) Yes, if $\sigma$ is large enough.
Univariate $m_{3}$ (Section 3.1) $\nabla m_{3}(s^{*})=0$ and $H+\frac{1}{3}T[s^{*}]+\sigma\|s^{*}\|^{2}\geq\frac{T^{2}}{18\sigma}$ Yes, for all $\sigma>0$.
QQR model ($T=0$) (Theorem 3.3) $\nabla m_{3}(s^{*})=0$ and $H+\sigma\|s^{*}\|^{2}I_{n}\succeq 0$ Yes, for all $\sigma>0$.
Special $T$ model $T[s]^{3}=t^{T}s\|s\|^{2}$ (Section 3.2) $\nabla m_{3}(s^{*})=0$ and $B_{T}(s^{*})\succeq 0$ (37) Yes for all $\sigma>0$.
Schnabel Model (Section 3.3) $\nabla m_{\mathcal{S}}(s^{*})=0$ and $B_{\mathcal{S}}(s^{*})\succeq 0$ (41) Sometimes no, even if $\sigma$ is large.
3.1 Univariate quartically regularized cubic mode $(n=1)$
The global optimality characterization of the univariate quartically regularized polynomial is given in Lemma 3.1.
Lemma 3.1.
For $n=1$, $s^{*}$ satisfies
$\nabla m_{3}(s^{*})=0\qquad\text{and}\qquad H+\frac{1}{3}T[s^{*}]+\sigma\|s^{*}\|^{2}\geq\frac{T^{2}}{18\sigma},$ (36)
if and only if $s^{*}$ is a global minimizer.
Proof.
Let $q(v):=m_{3}(s^{*}+v)-m_{3}(s^{*})$, using (5), we have
$2q(v)$ $=$ $2\underbrace{\nabla m_{3}(s^{*})^{T}}_{=0}v+\frac{1}{2}v^{T}\bigg[\nabla^{2}m_{3}(s^{*})[v]^{2}+\frac{1}{3}T[v]^{3}+2\sigma{s^{*}}^{T}v+\frac{\sigma}{2}\|v\|^{2}\bigg]v$
$\underset{\text{use }n=1}{=}$ $\frac{1}{2}v^{T}\bigg[\bigg(H+\frac{1}{3}T[s^{*}]+\sigma\|s^{*}\|^{2}\bigg)+2\sigma{s^{*}}^{2}+\frac{2}{3}T[s^{*}]+\frac{1}{3}T[v]+2\sigma{s^{*}}v+\frac{\sigma}{2}v^{2}\bigg]v$
$=$ $\frac{1}{2}v^{T}\bigg[H+\frac{1}{3}T[s^{*}]+\sigma\|s^{*}\|^{2}-\frac{T^{2}}{18\sigma}\bigg]v+\frac{\sigma}{4}\bigg(v+2s^{*}-\frac{T}{3\sigma}\bigg)^{2}v^{2}.$
$\Rightarrow$’ Clearly, if (36) is satisfied, $m_{3}(s^{*}+v)-m_{3}(s^{*})\geq 0$ for all $v$. Therefore, $s^{*}$ is the global minimum.
$\Leftarrow$’ If $s^{*}$ is a global minimum, then it is a stationary point, therefore $\nabla m_{3}(s^{*})=0$. Since $s^{*}$ is the global minimum $m_{3}(s^{*}+v)-m_{3}(s^{*})\geq 0$ for all $v\in\mathbb{R}^{n}$. By setting $v=-(2s^{*}+\frac{1}{3\sigma}T)$, we derive that $H+\frac{1}{3}T[s^{*}]+\sigma\|s^{*}\|^{2}\geq\frac{T^{2}}{18\sigma}$. ∎
Hilbert [30] proved that every nonnegative univariate polynomial is SoS. Here we provide an alternative proof of this result using global optimality conditions.
Theorem 3.1.
For $n=1$, if $s^{*}$ is a global minimizer, then $q(v)=m_{3}(s^{*}+v)-m_{3}(s^{*})$ is SoS.
Proof.
Since $s^{*}$ is the global minimum, we have $\nabla m_{3}(s^{*})=0$, according to (5), we can write
$m_{3}(s^{*}+v)-m_{3}(s^{*})=\frac{1}{2}\begin{bmatrix}1\\ v\\ v^{2}\end{bmatrix}\underbrace{\begin{bmatrix}0&0&0\\ 0&\nabla^{2}m_{3}(s^{*})&\sigma s^{*}+\frac{T}{6}\\ 0&\sigma s^{*}+\frac{T}{6}&\frac{\sigma}{2}\end{bmatrix}}_{Q_{a}}\begin{bmatrix}1\\ v\\ v^{2}\end{bmatrix}$.
Since $s^{*}$ is the global minimizer, by Lemma 3.1, (36) holds and
[…]
Note that the term in $3\sigma\|s^{*}\|^{2}$ is only true for $n=1$ case, in $n\geq 2$, the term will be $\sigma(\|s^{*}\|^{2}I_{n}+ss^{T})$. Therefore, The determinant of the submatrix in $Q_{a}$,
$\frac{\sigma}{2}\nabla^{2}m_{3}(s^{*})-\bigg(\sigma s^{*}+\frac{T}{6}\bigg)^{2}\geq\sigma^{2}\|s^{*}\|^{2}+\frac{\sigma}{3}Ts^{*}+\frac{T^{2}}{36}-\bigg(\sigma s^{*}+\frac{T}{6}\bigg)^{2}=0.$
We deduce that $Q_{a}\succeq 0$, therefore $m_{3}(s^{*}+v)-m_{3}(s^{*})$ is SoS. ∎
Remark 3.1.
When $T=0$, (36) becomes $\nabla m_{3}(s^{*})=0$ and $H+\sigma\|s^{*}\|^{2}I_{n}\succeq 0$ which coincides with the necessary and sufficient global optimality condition for the QQR polynomial in (16).
Remark 3.2.
Consider global optimality condition in [76, Thm 2.1 and 2.2], for $n=1$, (Necessary cond.) gives to $H+\frac{1}{3}T[s^{*}]+\sigma\|s^{*}\|^{2}\geq-\frac{1}{3}T[s^{*}]+\frac{|T|}{3}\|s^{*}\|\geq 0$ and (Sufficient cond.) gives to $H+\frac{1}{3}T[s^{*}]+\sigma\|s^{*}\|^{2}\geq\frac{{\hat{T}}^{2}}{18\sigma}-\frac{1}{3}T[s^{*}]+\frac{|T|}{3}\geq\frac{{\hat{T}}^{2}}{18\sigma}.$ note that (36) lies between (Necessary cond.) and (Sufficient cond.), such that
[…]
Lemma 3.1 provides a tighter necessary and sufficient condition for global optimality in the $n=1$ case.
3.2 Quartically Regularized Cubic Model with Special Tensor Term $T$
We can also obtain tight necessary and sufficient global optimality conditions for the quartically regularized problem when the cubic tensor has the special structure $T[s]^{3}=t^{T}s\,\|s\|^{2}$.
Lemma 3.2.
If for any $s\in\mathbb{R}^{n}$, $T$ satisfies $T[s]^{3}=t^{T}s\|s\|^{2}$ for some $t\in\mathbb{R}^{n}$ then $s^{*}$ satisfies
$\nabla m_{3}(s^{*})=0\qquad\text{and}\qquad B_{T}(s^{*}):=H+T[s^{*}]+\sigma\|s^{*}\|^{2}I_{n}-\frac{1}{3}({s^{*}}t^{T}+t{s^{*}}^{T})-\frac{tt^{T}}{18\sigma}\succeq 0,$ (37)
if and only if $s^{*}$ is a global minimizer of $m_{3}$.
Proof.
Let $q(v):=m_{3}(s^{*}+v)-m_{3}(s^{*})$. Using the expression of $B_{T}(s^{*})$, we have
$\nabla^{2}m_{3}(s^{*})=B_{T}(s^{*})+2\sigma s^{*}{s^{*}}^{T}+\frac{1}{3}({s^{*}}t^{T}+t{s^{*}}^{T})+\frac{tt^{T}}{18\sigma}I_{n}$
Substituting into (5), we have
$2\big[q(v)-\nabla m_{3}(s^{*})^{T}v\big]$ $=$ $v^{T}\bigg[\underbrace{B_{T}(s^{*})+2\sigma s^{*}{s^{*}}^{T}+\frac{1}{3}({s^{*}}t^{T}+t{s^{*}}^{T})+\frac{tt^{T}}{18\sigma}I_{n}}_{=\nabla^{2}m_{3}(s^{*})}+\frac{1}{3}T[v]+2\sigma{s^{*}}^{T}v+\frac{\sigma}{2}\|v\|^{2}I_{n}\bigg]v$
$=$ $B_{T}(s^{*})[v]^{2}+2\sigma(v^{T}s^{*})^{2}+\frac{2}{3}(v^{T}{s^{*}})(t^{T}v)+\frac{(v^{T}t)^{2}}{18\sigma}+\frac{1}{3}(t^{T}v)\|v\|^{2}+2\sigma{s^{*}}^{T}v\|v\|^{2}+\frac{\sigma}{2}\|v\|^{4}$
$=$ $B_{T}(s^{*})[v]^{2}+\frac{\sigma}{2}\bigg(\|v\|^{2}+2v^{T}s^{*}+\frac{v^{T}t}{3\sigma}\bigg)^{2}$
$=$ $B_{T}(s^{*})[v]^{2}+\frac{\sigma}{2}\left[\bigg\|v+\big(s^{*}+\frac{t}{6\sigma}\big)\bigg\|^{2}-\bigg\|s^{*}+\frac{t}{6\sigma}\bigg\|^{2}\right]^{2}$
$\Rightarrow$’ Clearly, if (37) is satisfied, $m_{3}(s^{*}+v)-m_{3}(s^{*})\geq 0$ for all $v$. Therefore, $s^{*}$ is the global minimum.
$\Leftarrow$’ If $s^{*}$ is a global minimum, then it is a stationary point, therefore $\nabla m_{3}(s^{*})=0$. Let $w:=v+2s^{*}+\frac{t}{3\sigma}$. Assume $w^{T}v=0$, since $s^{*}$ is the global minimum, we have $m_{3}(s^{*}+v)-m_{3}(s^{*})\geq 0$ for all $v\in\mathbb{R}^{n}$, thus $B_{T}(s^{*})[v]^{2}\geq 0$ for $v$ orthogonal to $w$.
Otherwise, assume $w:=v+2s^{*}+\frac{t}{3\sigma}$ and $s_{c}:=s^{*}+\frac{t}{6\sigma}$ are not orthogonal, the line $s_{c}+\tilde{k}w$ intersects the ball centred at the origin of radius $\|s_{c}\|$ at two points, $s_{c}$ and $u_{*}$, with $\|u_{*}\|=\|s_{c}\|.$ We set $v=-s_{c}+u_{*}=-\big(s^{*}+\frac{t}{6\sigma}\big)+u_{*}$. Since $s^{*}$ is the global minimum, we have $m_{3}(s^{*}+v)-m_{3}(s^{*})\geq 0$ for all $v\in\mathbb{R}^{n}$, we arrive at $B_{T}(s^{*})[v]^{2}\geq 0$ for $v$ not orthogonal to $w$. ∎
Theorem 3.2.
For $n\geq 2$, assume for any $s\in\mathbb{R}^{n}$, $T$ satisfies $T[s]^{3}=t^{T}s\|s\|^{2}$ for some $t\in\mathbb{R}^{n}$. If $s^{*}$ is the global minimizer, then $q(v)=m_{3}(s^{*}+v)-m_{3}(s^{*})$ is SoS.
Proof.
Since $s^{*}$ is the global minimum, we have $\nabla m_{3}(s^{*})=0$, according to (5), we can write
$m_{3}(s^{*}+v)-m_{3}(s^{*})=\frac{1}{2}\omega\underbrace{\begin{bmatrix}0&0&0&0&\\ 0&\nabla^{2}m_{3}(s^{*})&(\sigma s^{*}+\frac{t}{6})\mathbf{1}_{n}^{T}&0\\ 0&\mathbf{1}_{n}(\sigma s^{*}+\frac{t}{6})^{T}&\frac{\sigma}{2}\mathbf{1}_{n}\mathbf{1}_{n}^{T}&0\\ 0&0&0&0\end{bmatrix}}_{Q_{c}}\omega$ (38)
where $\phi_{2}$ defined as in (2), $\mathbf{1}=[1,\dotsc,1]^{T}\in\mathbb{R}^{n}$ and $\mathbf{1}\mathbf{1}^{T}\in\mathbb{R}^{n\times n}$ is a matrix with all entries equal to 1s. Let $\omega\in\mathbb{R}^{\frac{1}{2}(n+2)(n+1)}$ be any vector, we write $\omega=[1,v,u,z]^{T}$ as defined in (3). For any $\omega$, or equivalently for any $(u,v,z)$, we have
$2\omega^{T}Q_{c}(v)\omega$ $=$ $\nabla^{2}m_{3}(s^{*})[v]^{2}+2v^{T}(\sigma s^{*}+\frac{t}{6})\sum_{i=1}^{n}u_{i}+\frac{\sigma}{2}\sum_{i=1}^{n}u_{i}^{2}+\sigma\sum_{1\leq i<j\leq n}u_{i}u_{j}$ (39)
$=$ $\nabla^{2}m_{3}(s^{*})[v]^{2}+2v^{T}(\sigma s^{*}+\frac{t}{6})e^{T}u+\frac{\sigma}{2}(e^{T}u)^{2}$
$=$ $\underbrace{\nabla^{2}m_{3}(s^{*})[v]^{2}-2\sigma\bigg[v^{T}(s^{*}+\frac{t}{6\sigma})\bigg]^{2}}_{:\mathcal{J}_{1}}+\frac{\sigma}{2}\bigg[\sum_{i=1}^{n}u_{i}+2v^{T}(s^{*}+\frac{t}{6\sigma})\bigg]^{2}.$
Since $s^{*}$ is the global minimum, Lemma 3.2 proves that (37) holds. Rearranging (37) gives $\nabla^{2}m_{3}(s^{*})[v]^{2}\geq\frac{1}{3}({s^{*}}t^{T}+t{s^{*}}^{T})+\frac{tt^{T}}{18\sigma}+2\sigma s^{*}{s^{*}}^{T}$ for all $v\in\mathbb{R}^{n}$. Consequently,
$2\omega^{T}Q_{c}(v)\omega\geq\mathcal{J}_{1}\geq v^{T}\bigg[\frac{1}{3}({s^{*}}t^{T}+t{s^{*}}^{T})+\frac{tt^{T}}{18\sigma}I_{n}+2\sigma s^{*}{s^{*}}^{T}\bigg]v-2\sigma v^{T}\bigg[(s^{*}+\frac{t}{6\sigma})(s^{*}+\frac{t}{6\sigma})^{T}\bigg]v\geq 0$
for all $v\in\mathbb{R}^{n}$. Thus, we obtained that $Q_{c}\succeq 0$ and therefore $m_{3}(s^{*}+v)-m_{3}(s^{*})$ is SoS. ∎
Remark 3.3.
(Extending to $W$-norm) By changing $I_{n}$ to $W$-norm with $W\succeq 0$, if the model is bounded below or $W\succ 0$, then Theorem 3.2 and Lemma 3.2 can be extended to allow $T[s]^{3}=(t^{T}s)(s^{T}Ws)$ by using the change of $\tilde{s}=\hat{D}^{1/2}\hat{L}s$ where $\hat{L}^{T}\hat{D}\hat{L}=W$.
Quadratic quartically regularized model.
By setting $t=0$, Lemma 3.2 can be simplified to Lemma 3.3, which coincides with the result in [16, Theorem 8.2.8]. Also, Theorem 3.2 can be simplified to Theorem 3.3.
Lemma 3.3.
$s^{*}$ satisfies $\nabla m_{3}(s^{*})=0$ and $B_{QQR}:=H+\sigma\|s^{*}\|^{2}I_{n}\geq 0$ if and only if $s^{*}$ is a global minimizer.
Theorem 3.3.
For $n\geq 2$, if $T=0$, $s^{*}$ is the global minimum of $m_{3}$, then $q(v)=m_{3}(s^{*}+v)-m_{3}(s^{*})$ is SoS. Proof can be found in Section B.1
A discussion for $n=2$ case.
Hilbert proved that every nonnegative bivariate quartic polynomial is SoS [30]. Therefore, when $n=2$, the polynomial $q(v)=m_{3}(s^{*}+v)-m_{3}(s^{*})$ is SoS. Notice that in this case $T[v]^{3}=t_{1}v_{1}^{3}+t_{2}v_{1}^{2}v_{2}+t_{3}v_{1}v_{2}^{2}+t_{4}v_{2}^{3},$ and there are no cross terms of the form $t_{i,j,k}v_{i}v_{j}v_{k}$ with distinct indices $(i,j,k)$. This structure resembles the simplified tensor form $t^{T}v\,\|v\|^{2}$ in Theorem 3.2. We further conjecture that, for $n=2$, a necessary and sufficient global optimality condition (with no relaxation gap) holds for any tensor $T$. If this conjecture is true, the result may extend to any dimension $n\geq 2$ in which the tensor term $T[v]^{3}$ contains no cross terms $v_{i}v_{j}v_{k}$ with all distinct $(i,j,k)$, yielding corresponding global optimality certificates. We leave this investigation to future work. A detailed proof showing that every nonnegative bivariate quartic polynomial is SoS is due to Hilbert’s classical 1888 result on binary quartic forms [30, 64, 61], and this proof does not use the optimality conditions for the quartically regularized subproblem and is universal for any nonnegative bivariate quartic polynomial. The proof for binary quartic forms [30, 64, 61] cannot be directly generalized to higher dimensions due to insufficient degrees of freedom for cubic cancellation. For $n=3$, a homogeneous cubic terms involves $\binom{3+3-1}{3}=10$ monomial coefficients, whereas the expression $t^{T}s\|s\|^{2}$ (or $t^{T}s\|s\|_{W}^{2}$) where $t^{T}s$ is linear and $\|s\|$ (or $\|s\|_{W}^{2}$) is quadratic, has only three to six independent coefficients. Hence, in general one cannot find $t^{T}s\|s\|^{2}$ that exactly reproduces an arbitrary homogeneous cubic terms. The lack in degree of freedom prevents the complete elimination of the cubic component.
3.3 The Schnabel Model
Schnabel et al. [66] proposed a formulation for unconstrained optimization based on a fourth-order model of the objective function $f$. The model is given by
$m_{\mathcal{S}}(s):=f_{0}+g^{T}s+\frac{1}{2}H[s]^{2}+\frac{1}{6}\sum_{j=1}^{k}(a_{j}^{T}s)^{2}(b_{j}^{T}s)+\frac{1}{4}\sum_{j=1}^{k}{\sigma_{j}}(a_{j}^{T}s)^{4}$ (Schnabel Model)
where $\sigma_{j}>0$, $a_{j},b_{j}\in\mathbb{R}^{n}$ are specific vectors which correspond to the low–rank approximations of the third and fourth-order derivatives and are built by interpolating function and gradient information from past iterates. Schnabel et al. showed that although (Schnabel Model) is a fourth-order polynomial in $n$ variables, it can be reduced to the minimization of a fourth-order polynomial in $k$ variables plus a quadratic in nkn-k variables. Note that the quartic term in $m_{\mathcal{S}}$ does not necessarily ensure that $m_{\mathcal{S}}(s)\to+\infty$ as $\|s\|\to\infty$ in all directions. Therefore, the Schnabel model may not be bounded below and cannot guarantee the existence of a global minimizer of the model function. In [66, 67], Schnabel et al. used this model within a trust-region framework to obtain global convergence of the overall algorithm.
Let $q_{\mathcal{S}}(v):=m_{\mathcal{S}}(s+v)-m_{\mathcal{S}}(s)$. We begin with a technical lemma that provides alternative equivalent expressions for $q_{\mathcal{S}}(v)$. This lemma is a simplified version of the result in Lemma 2.1.
Lemma 3.4.
For any vector $s,v\in\mathbb{R}^{n}$, let $q_{\mathcal{S}}(v):=m_{\mathcal{S}}(s+v)-m_{\mathcal{S}}(s)$. We have
$q_{\mathcal{S}}(v):=\nabla m_{\mathcal{S}}(s)^{T}v+\frac{1}{2}\nabla^{2}m_{\mathcal{S}}(s)[v]^{2}+\sum_{j=1}^{k}\bigg[\sigma_{j}(a_{j}^{T}v)^{2}(s^{T}a_{j})+\frac{1}{6}(a_{j}^{T}v)^{2}(b_{j}^{T}v)+\frac{\sigma_{j}}{4}(a_{j}^{T}v)^{4}\bigg].$ (40)
where $\nabla^{2}m_{\mathcal{S}}(s)=H+\sum_{j=1}^{k}(a_{j}a_{j}^{T})(b_{j}^{T}s)+\sum_{j=1}^{k}3\sigma_{j}(a_{j}a_{j}^{T})(a_{j}^{T}s)^{2}.$
Previously, in Lemma 3.2, we established tight necessary and sufficient conditions that fully characterize the global minimizer for $m_{3}$. For (Schnabel Model), however, there is an inherent gap between necessary and sufficient conditions. Theorem 3.4 proves the sufficient conditions for global optimality and its relationship to SoS characterization. Theorem 3.5 proves the necessary conditions for global optimality
Theorem 3.4.
(SoS and Global Sufficient Optimality) For $n\geq 2$, assume $s^{*}$ satisfies
$\nabla m_{\mathcal{S}}(s^{*})=0\qquad\text{and}\qquad B_{\mathcal{S}}(s^{*}):=\nabla^{2}m_{\mathcal{S}}(s^{*})-2\sum_{i=1}^{k}\sigma_{j}\bigg[(a_{j}a_{j}^{T})s^{*}+\frac{b_{j}}{6\sigma_{j}}\bigg]\bigg[(a_{j}a_{j}^{T})s^{*}+\frac{b_{j}}{6\sigma_{j}}\bigg]^{T}\succeq 0,$ (41)
then $q_{\mathcal{S}}(v)=m_{\mathcal{S}}(s^{*}+v)-m_{\mathcal{S}}(s^{*})$ is SoS. Also, if $s^{*}$ satisfies (41), then, $s^{*}$ is the global minimum.
Proof.
Using Lemma 3.4, (40) and (41), we have
$2q_{\mathcal{S}}(v)$ $=$ $v^{T}\bigg[B_{\mathcal{S}}(s^{*})+2\sum_{i=1}^{k}\sigma_{j}\bigg((a_{j}a_{j}^{T})s^{*}+\frac{b_{j}}{6\sigma_{j}}\bigg)\bigg((a_{j}a_{j}^{T})s^{*}+\frac{b_{j}}{6\sigma_{j}}\bigg)^{T}\bigg]v$ (42)
$\qquad+2\sum_{i=1}^{k}\bigg[\sigma_{j}(a_{j}^{T}v)^{2}s^{T}(a_{j}a_{j}^{T})v++\frac{1}{2}(a_{j}^{T}v)^{2}(b_{j}^{T}v)+\frac{\sigma_{j}}{4}(a_{j}^{T}v)^{4}\bigg]$
$=$ $\underbrace{B_{\mathcal{S}}(s^{*})}_{\succeq 0}[v]^{2}+\sum_{j=1}^{k}\frac{\sigma_{j}}{2}\bigg[(a_{j}^{T}v)^{2}+2(a_{j}^{T}s^{*})(a_{j}^{T}v)+\frac{b_{j}^{T}v}{3\sigma_{j}}\bigg]^{2}.$
Note that the deduction of (42) is similar to the proof of Lemma 3.1. Since (42) is symmetric and positive semidefinite, the quadratic form $B_{\mathcal{S}}[v]^{2}$ is a sum of squares. Together with the remaining sum-of-squares terms, this implies that $q_{\mathcal{S}}(v)$ is itself a sum-of-squares polynomial. To prove the second statement, $q_{\mathcal{S}}(v):=m_{\mathcal{S}}(s^{*}+v)-m_{\mathcal{S}}(s^{*})$ is SoS and thus nonnegative for all $v$. Therefore, $s^{*}$ is the global minimum. ∎
Remark 3.4.
Using the expression of $\nabla^{2}m_{\mathcal{S}}$ in Lemma 3.4, the alternative formulation of $B_{\mathcal{S}}(s^{*})$ is
$B_{\mathcal{S}}(s):=H+T[s]+\sum_{j=1}^{k}\sigma_{j}(a_{j}a_{j}^{T})(a_{j}^{T}s)^{2}-\sum_{j=1}^{k}\frac{a_{j}a_{j}^{T}}{3}\bigg[sb_{j}^{T}+b_{j}{s}^{T}\bigg]-\sum_{j=1}^{k}\frac{b_{j}b_{j}^{T}}{18\sigma_{j}}$ (43)
where $T=\sum_{j=1}^{k}a_{j}\circ a_{j}\circ b_{j}$. This condition can be viewed as a generalization of (37). From Theorem 3.1Theorem 3.4, we observe that the SoS conditions are derived in more general forms.
Remark 3.5.
(Separable Regularization Norm) To ensure that (Schnabel Model) is bounded below. Set $k=n$ and, without loss of generality, assume that $\sigma_{j}=\sigma$ and $a_{j}=e_{j}$, where $e_{j}$ denotes the unit vector with 1 in the $j$-th entry and 0 elsewhere. Note that any other choice of $\sigma_{j},a_{j}$ can be reduced to this form through an appropriate change of variables. Let $T=\sum_{j=1}^{n}e_{j}\circ e_{j}\circ b_{j}$, (Schnabel Model) simplifies to
$m_{\mathcal{S}}(s)$ $=f_{0}+g^{T}s+\frac{1}{2}H[s]^{2}+\frac{1}{6}T[s]^{3}+\frac{\sigma}{4}\|s\|_{4}^{4},\qquad\text{where}\qquad\|s\|_{4}=\big(\sum_{j=1}^{n}s_{j}^{4}\big)^{1/4},$
The key difference between this model and the quartically regularized polynomial model $m_{3}$ lies in the regularization term: here we employ a quartically separable norm $\|s\|_{4}^{4}=\sum_{j=1}^{n}s_{j}^{4}$, whereas $m_{3}$ uses a quartically regularized Euclidean norm $\|s\|_{2}^{4}=(s_{1}^{2}+\dotsc s_{n}^{2})^{2}$. Note that under this construction, $B_{\mathcal{S}}(s)$ has the expression,
$B_{\mathcal{S}}(s):=H+T[s]+\sigma\,\mathrm{diag}\{s_{j}^{2}\}_{1\leq j\leq n}-\sum_{j=1}^{n}\frac{e_{j}e_{j}^{T}}{3}\big(sb_{j}^{T}+b_{j}s^{T}\big)-\sum_{j=1}^{n}\frac{b_{j}b_{j}^{T}}{18\sigma_{j}}.$ (44)
Theorem 3.5.
(Global Necessary Optimality) Let $\text{Im}(A):=\operatorname{span}\{a_{1},\dotsc,a_{k}\}$ and its orthogonal complement space $\text{Im}(A)^{\perp}:=\operatorname{span}\{a_{k+1},\dotsc,a_{n}\}$. If a global minimizer $s^{*}$ exists and is attained, then $\nabla m_{\mathcal{S}}(s^{*})=0$ and
$B_{\mathcal{S}}(s^{*})[v]^{2}\geq 0\qquad\text{for all}\qquad v\in\text{Im}(A)^{\perp}.$
If $\{a_{1},\dotsc,a_{k}\}$ are orthogonal vectors, then we also have
$B_{\mathcal{S}}(s^{*})[a_{j}]^{2}\geq 0\qquad\text{for all}\qquad a_{j}\in\text{Im}(A).$
Proof.
If a global minimum exists and $s^{*}$ is a global minimum, then it is a stationary point; therefore $\nabla m_{\mathcal{S}}(s^{*})=0$. From (42),
$2q_{\mathcal{S}}(v)=B_{\mathcal{S}}(s^{*})[v]^{2}+\sum_{j=1}^{k}\frac{\sigma_{j}}{2}\left[\bigg\|a_{j}a_{j}^{T}v+a_{j}a_{j}^{T}s^{*}+\frac{b_{j}}{6\sigma_{j}}\bigg\|^{2}-\bigg\|a_{j}a_{j}^{T}s^{*}+\frac{b_{j}}{6\sigma_{j}}\bigg\|^{2}\right]^{2}.$ (45)
Note that the deduction of (45) is similar to is similar to the proof of Lemma 3.1.
  1. 1.
    If $v\in\text{Im}(A)^{\perp}$, then $a_{j}a_{j}^{T}v=0$ for all $j=1,\dotsc,k$. Then, we have $B_{\mathcal{S}}(s^{*})[v]^{2}\geq 0$ for $v\in\text{Im}(A)^{\perp}$.
  2. 2.
    Otherwise, for each unit vector $v=\tilde{k}\frac{a_{j}}{\|a_{j}\|}\in\text{Im}(A)$, since $a_{j}^{T}a_{\iota}=0$ for $\iota\neq 0$. Therefore, let $s_{c}:=a_{j}a_{j}^{T}s^{*}+\frac{b_{j}}{6\sigma_{j}}$,
    $2q_{\mathcal{S}}(v)=\tilde{k}^{2}\|a_{j}\|^{-2}B_{\mathcal{S}}(s^{*})[a_{j}]^{2}+\frac{\sigma_{j}}{2}\left[\big\|\tilde{k}a_{j}+s_{c}\big\|^{2}-\big\|s_{c}\big\|^{2}\right]^{2}.$ (46)
[…]
A natural question arises: is $B_{\mathcal{S}}(s^{*})$ positive semidefinite when $\sigma$ is sufficiently large? Equivalently, if $\sigma$ is large enough, can $m_{\mathcal{S}}(s)-m_{\mathcal{S}}(s^{*})$ be represented as a SoS? Unfortunately, unlike the case of $m_{3}$, the answer is no. An illustrative counterexample is provided in Section 4.1. If we attempt to extend the analysis from Theorem 2.2 and Theorem 2.3 to the current setting, we find that the bound $\|s^{*}\|=O(\|g\|^{1/3})$ in Theorem 2.2 still holds—with minor adjustments replacing $\|s^{*}\|$ by $\|s^{*}\|_{4}$. However, difficulties arise in the proof of (31) in Theorem 2.3. In $m_{3}$, the regularization term of $B(s^{*})$ is $\sigma\|s^{*}\|^{2}I_{n}$, whose minimum eigenvalue equals $\sigma\|s^{*}\|^{2}$. In contrast, the regularization term in $B_{\mathcal{S}}(s^{*})$ is $\sigma\,\mathrm{diag}\{s_{j}^{2}\}_{1\leq j\leq n}$ as given in (44), whose minimum eigenvalue is $\sigma\min_{1\leq j\leq n}\{s_{j}^{2}\}$. The step-size bound $\|s^{*}\|=O(\epsilon_{g}^{1/3})$ does not directly control this minimum component $\min_{j}|s_{j}|$, and hence does not guarantee that the regularization term in $B_{\mathcal{S}}(s^{*})$ dominates the remaining terms. Consequently, $B_{\mathcal{S}}(s^{*})$ may fail to be PSD. This demonstrates that even a subtle change in the form of the regularization norm can fundamentally alter the SoS properties of the model as $\sigma\to\infty$.
Difference between the special $T$ model and the Schnabel model
Under special cases, (Schnabel Model) can be transformed into our special $T$ with a $W$ semi-norm ($W\succeq 0$). If $b_{1}=b_{2}=\dotsc=b_{k}$ and $\sigma_{1}=\sigma_{2}=\dotsc=\sigma_{k}=\frac{3}{2}\sigma$, we can set $b_{1}:=\frac{t}{3}$ and $\sigma_{1}:=\sigma$,
$m_{\mathcal{S}}(s):=f_{0}+g^{T}s+\frac{1}{2}H[s]^{2}+\frac{1}{6}\bigg[s^{T}\underbrace{\big(\sum_{j=1}^{k}a_{j}a_{j}^{T}\big)}_{:=W}s\bigg](t^{T}s)+\frac{\sigma}{4}\|s\|_{W}^{4}.$ (47)
Therefore, we have $T[s]^{3}=(t^{T}s)(s^{T}Ws)$ and clearly $W\succeq 0$. Under this specific constructions, if the model is bounded below, then $q_{\mathcal{S}}(v)=m_{\mathcal{S}}(s^{*}+v)-m_{\mathcal{S}}(s^{*})$ is SoS (see Lemma 3.2 and LABEL:{extending_to_W_norm}).
Remark 3.6.
(Difference between the regularization norm) There is no matrix transform / linear change of basis that converts $\sum_{j=1}^{k}s_{j}^{4}$ into the Euclidean norm $\|s\|^{4}=\big(\sum_{j=1}^{k}s_{j}^{2}\big)^{2}$ for $j\geq 2$. More generally, the regularization term in (Schnabel Model) takes the form $\sum_{j=1}^{k}(a_{j}^{T}s)^{4}=\sum_{j=1}^{k}(s^{T}a_{j}a_{j}^{T}s)^{2}=\sum_{j=1}^{k}\|s\|_{W_{j}}^{4},\qquad W_{j}=a_{j}a_{j}^{T}.$ Here, $\|s\|_{W_{j}}:=\sqrt{s^{T}W_{j}s}$ is the matrix-induced quadratic form. Since each $W_{j}$ is rank-one and positive semidefinite, $\|\cdot\|_{W_{j}}$ is not a norm on the whole space $\mathbb{R}^{n}$, but only a seminorm: it vanishes for any $s\perp a_{j}$, even when $s\neq 0$. Only in the case $k=1$ does the regularization reduce to a single quartic seminorm, i.e.,
$(a_{1}^{T}s)^{4}=\|s\|_{W_{1}}^{4},$
which is in a form analogous to the quartic regularization norm in the special $T$ model (i.e., $\|s\|_{W}^{4}$ or $\|s\|^{4}$). The quartic semi-norm does not necessarily ensure that $m_{\mathcal{S}}$ is bounded below. For instance, if $s\in\operatorname{Im}(A)^{\perp}$ and $H[s]^{2}<0$, then $m_{\mathcal{S}}(s)\to$-$\infty$ as $\|s\|\to\infty$.
Remark 3.7.
When $k=1$, $m_{\mathcal{S}}(s):=f_{0}+g^{T}s+\frac{1}{2}H[s]^{2}+\frac{1}{6}(a_{1}^{T}s)^{2}(b_{1}^{T}s)+\frac{\sigma_{1}}{4}(a_{1}^{T}s)^{4}.$ In this case, the global necessary optimality condition in Theorem 3.5 becomes $\nabla m_{\mathcal{S}}(s^{*})=0$ and $B_{\mathcal{S}}(s^{*})\succeq 0$. Hence, if a global minimizer exists, the necessary and sufficient conditions coincide and $m_{\mathcal{S}}(s)-m_{\mathcal{S}}(s^{*})$ is SoS. This corresponds to the special case in (47), which can be transformed into the special $T$-model with a weighted quartic semi-norm. For $k>1$, however, the presence of multiple distinct seminorms $\|s\|_{W_{j}}$ prevents such a change of basis and reduction; the equivalence no longer holds.
4 Examples and Preliminary Numerical Results
In this section, we begin by examining several well-known quartic examples due to Nesterov and Ahmadi (Section 4.1 and Section 4.2). These problems are known to be NP-hard to globally minimize and are not SoS-exact at the zeroth level of the Lasserre hierarchy. We compare these examples with their quartically regularized counterparts and analyze—both theoretically and numerically—how different regularization norms and problem constraints affect SoS representability.
In Section 4.3, we present a third-order algorithm that leverages the SoS property of sufficiently regularized nonnegative quartic polynomials. The algorithm is built upon the AR33 framework; at each iteration, the regularization parameter is chosen to satisfy two requirements: (1) guaranteeing a sufficient decrease in the objective value, and (2) ensuring that the subproblem becomes globally solvable in polynomial time. We test this method on the Moré–Garbow–Hillstrom (MGH) test set [42] and provide preliminary numerical observations on several representative and challenging examples, including the Nesterov–Chebyshev function, the Beale function, and the sensor network localization problem.
4.1 Examples with Alternative Quartic Norms
In the special $T$-model or the QQR model, which uses a quartic Euclidean norm, $\|s\|^{4}\;=\;\big(\sum_{i=1}^{n}s_{i}^{2}\big)^{2},$ the shifted polynomial $m_{3}(s)-m_{3}(s^{*})$ is SoS for all $\sigma>0$. However, if we replace the Euclidean norm with the quartic $\ell_{4}$-norm, $\|s\|_{4}^{4}\;=\;\sum_{i=1}^{n}s_{i}^{4},$ then the shifted polynomial need not remain SoS for all $\sigma>0$. We illustrate using a modified example inspired by Ahmadi et al. in [2, Thm. 3.3]. Consider a trivariate quartic polynomial,
$m_{\mathcal{A}}(s)=2\|s\|^{2}+8(s_{1}s_{2}+s_{1}s_{3}+s_{2}s_{3})+\frac{\sigma}{4}(s_{1}^{4}+s_{2}^{4}+s_{3}^{4})=\hat{H}[s]^{2}+\frac{\sigma}{4}\sum_{i=1}^{3}s_{i}^{4},\qquad\sigma>0.$ (48)
where $\hat{H}=[4,8,8;8,4,8;8,8,4]\in\mathbb{R}^{3\times 3}$. $\nabla^{2}m_{\mathcal{A}}(0)=\hat{H}$ which has eigenvalue [4,4,20][-4,-4,20]. Clearly, $m_{\mathcal{A}}(s)$ is nonconvex for all $\sigma>0$.
Lemma 4.1.
For any $\sigma>0$, let $m_{\mathcal{A}}(s)$ be defined as in (48). Denote its global minimum value by $m_{\mathcal{A}}(s^{*})$. Then, for every $\sigma>0$, the polynomial $m_{\mathcal{A}}(s)-m_{\mathcal{A}}(s^{*})$ is nonnegative but not a SoS.
Proof.
For the specific case of $\sigma=4$, we denote
$m_{\mathcal{A},\sigma_{0}}(s):=2\|s\|^{2}+8(s_{1}s_{2}+s_{1}s_{3}+s_{2}s_{3})+(s_{1}^{4}+s_{2}^{4}+s_{3}^{4})=\hat{H}[s]^{2}+\sum_{i=1}^{3}s_{i}^{4}.$
Note that $m_{\mathcal{A},\sigma_{0}}(s)$ is bounded below and possesses six global minimizers [^4] Numerically, we obtain $m^{*}_{(\sigma_{0})}=-2.1443$, and $s^{*}_{(\sigma_{0})}=\pm(-1.1498,\,0.6674,\,0.6674)$ and their permutations.. We denote the global minimizers and global minimum as $(s^{*}_{(\sigma_{0})},m^{*}_{(\sigma_{0})})$. [2, Thm 3.3] proves that, the shifted polynomial $m_{\mathcal{A},\sigma_{0}}(s)-m^{*}_{(\sigma_{0})}$ is not SoS in $s$. We prove that this result extends to all $\sigma>0$. Take any fixed $\sigma>0$, consider the change of variables $\tilde{s}=\frac{\sigma^{1/2}}{2}s$. Then, from (48), we have
[…]
By this scaling, $m_{\mathcal{A}}(s)$ attains the global minimizers and global minimum at $(2\sigma^{-1/2}s^{*}_{(\sigma_{0})},4\sigma^{-1}m^{*}_{(\sigma_{0})})$. As scaling preserve non-SoSness, given $m_{\mathcal{A},\sigma_{0}}(\tilde{s})-m^{*}_{(\sigma_{0})}$ is not SoS, we deduce that $4\sigma^{-1}\big[m_{\mathcal{A},\sigma_{0}}(\tilde{s})-m^{*}_{(\sigma_{0})}\big]$ is also not SoS in $s$. Consequently, $m_{\mathcal{A}}(s)-m_{\mathcal{A}}(s^{*})=m_{\mathcal{A}}(s)-4\sigma^{-1}m^{*}_{(\sigma_{0})}$ is not SoS for any $\sigma>0$. ∎
Lemma 4.1 can be verified numerically. Using an SoS relaxation (e.g., SoSTOOLS) with $\sigma=4$, the moment-SDP solver returns a lower bound but not a global minimum, indicating that the SoS relaxation fails to recover the true optimum [^5] For very large $\sigma$ (e.g., $\sigma\geq 2500$), the solver reports a global minimum close to the origin, but this is due to numerical errors, not because the shifted polynomial becomes SoS..
Connection to the Schnabel model. For any $\sigma>0$, the polynomial (48) can be written in the form (Schnabel Model) by choosing $a_{j}=e_{j},b_{j}=0,$ for $j=1,2,3$ $g=0$, and $H=\hat{H}$. Using $s^{*}=2\sigma^{-1/2}s^{*}_{(\sigma_{0})}$ and $s^{*}_{(\sigma_{0})}=[-1.1498,\,0.6674,\,0.6674]^{T}$. We derive from (43), $B_{\mathcal{S}}(s^{*})$ for $m_{\mathcal{A}}$ has the expression
$B_{\mathcal{S}}(s^{*})=H+\operatorname{diag}({s^{*}_{(\sigma_{0})}}^{2})=\begin{bmatrix}9.2878&8&8\\ 8&5.7819&8\\ 8&8&5.7815\end{bmatrix}$ (49)
for any $\sigma>0$. $B_{\mathcal{S}}(s^{*})$ has eigenvalues [2.2,0,23.1][-2.2,0,23.1] and is therefore indefinite for all $\sigma>0$. Thus, the sufficient SoS condition in Theorem 3.4 fails. Yet, we verify that $s^{*}$, the necessary global condition in Theorem 3.5 is satisfied: $B_{\mathcal{S}}(s^{*})[e_{j}]^{2}\geq 0,j=1,2,3.$
Replacing the separable quartic norm. Let us now consider the Euclidean quartic regularization,
$m_{\mathcal{E}}(s)=2\|s\|^{2}+8(s_{1}s_{2}+s_{1}s_{3}+s_{2}s_{3})+\sigma\|s\|^{4}=\hat{H}[s]^{2}+\sigma\|s\|^{4}\qquad\sigma\geq 0.$ (50)
$m_{\mathcal{E}}(s)$ is a nonconvex quadratic quartically regularized (QQR) polynomial. By Theorem 3.3, we know that $m_{\mathcal{E}}(s)-m_{\mathcal{E}}(s^{*})$ is SoS for every $\sigma\geq 0$. In Lemma 4.2, we also derived the explicit expression for its global minimizers.
Lemma 4.2.
Let $s^{*}\in\mathbb{R}^{3}$ satisfy $s_{1}^{*}+s_{2}^{*}+s_{3}^{*}=0,\|s^{*}\|=\sigma^{-1/2}.$ Then $s^{*}$ is a global minimizer of $m_{\mathcal{E}}$, and its minimum value is $m_{\mathcal{E}}(s^{*})=-\sigma^{-1}.$
Proof.
$m_{\mathcal{E}}(s)+\sigma=\sigma\big(\|s\|^{2}-\sigma^{-1}\big)^{2}+4(s_{1}+s_{2}+s_{3})^{2},$ which is nonnegative and SoS for all $s$. The lower bound $\sigma^{-1}$ is attained precisely when $s_{1}+s_{2}+s_{3}=0$ and $\|s\|=\sigma^{-1/2}$. ∎
Lemma 4.2 shows that $m_{\mathcal{E}}$ has infinitely many global minimizers: every point on the circle $\{s\in\mathbb{R}^{3}:s_{1}+s_{2}+s_{3}=0,\ \|s\|=\sigma^{-1/2}\}.$ Although $m_{\mathcal{E}}$ is SoS, the moment relaxations numerical solvers (e.g., yalmip) return a rank-deficient moment matrix. This is because (50) has infinitely many global minimizers—the rank condition for finite relaxation cannot be satisfied. From (16), we have
$B_{QQR}({s^{*}}^{2})=H+\sigma\operatorname{diag}({s^{*}}^{2})=\begin{bmatrix}4+\sigma\|s^{*}\|^{2}&8&8\\ 8&4+\sigma\|s^{*}\|^{2}&8\\ 8&8&4+\sigma\|s^{*}\|^{2}\end{bmatrix}=\begin{bmatrix}8&8&8\\ 8&8&8\\ 8&8&8\end{bmatrix}$. (51)
$B_{QQR}({s^{*}}^{2})$ has eigenvalue $0,8,24$ and $B_{QQR}({s^{*}}^{2})\succeq 0$ for all $\sigma>0$. The subtle difference between (49) and (51) $B_{\mathcal{S}}(s^{*})$, the separable quartic norm does not guarantee positive definiteness, whereas the Euclidean quartic norm does.
Robustness: the effect of perturbations on global minimizers and on SoS representability. We introduce a small perturbation to (50), either by perturbing a single coefficient,
$\hat{m}_{\mathcal{E},1}(s)=\frac{\sigma}{4}\|s\|^{4}+2\|s\|^{2}+8\Big[(1+10^{-5})s_{1}s_{2}+s_{1}s_{3}+s_{2}s_{3}\Big].$
Alternatively, we can add small random perturbations to several entries
$\hat{m}_{\mathcal{E},2}(s)=\sigma\|s\|^{4}+2\|s\|^{2}+8\Big[(1+10^{-6}\,\xi_{12})s_{1}s_{2}+(1+10^{-6}\,\xi_{13})s_{1}s_{3}+(1+10^{-6}\,\xi_{23})s_{2}s_{3}\Big],$
where each $\xi_{ij}\sim\mathcal{N}(0,1)$ represents an independent Gaussian random variable. When we apply the moment-relaxation solver yalmip to both perturbed polynomials, it returns a successful status of 1, indicating that the ring of infinite global minimizers breaks—after perturbation—into one or at most two global minimizers. Consequently, the rank condition is satisfied. Small perturbations in the data (i.e., in the coefficients) can change the numerical outcome returned by yalmip. This behaviour observed arises due to the lack of robustness of this specific example.
Perturbing the cubic term. We also consider perturbing the cubic term (48),
$m_{\mathcal{A},2}(s)=2\|s\|^{2}+8(s_{1}s_{2}+s_{1}s_{3}+s_{2}s_{3})+0.1(s_{1}^{3}+s_{2}^{3}+s_{3}^{3})+\sigma(s_{1}^{4}+s_{2}^{4}+s_{3}^{4})$
for $\sigma=4$. This polynomial attains its global minimum at $s^{*}=[-1.1722,\ 0.6700,\ 0.6700]^{T},m_{\mathcal{A},2}(s^{*})=-2.2409.$ Using SoSTOOLS, we obtain a lower bound $\gamma^{*}=-2.3990$, which is strictly less than the true global minimum, implying that $m_{\mathcal{A},2}(s)-\gamma^{*}$ is still not SoS. In contrast, if we replace the separable quartic term by the Euclidean quartic term, we obtain the variant
$m_{\mathcal{E},2}(s)=2\|s\|^{2}+8(s_{1}s_{2}+s_{1}s_{3}+s_{2}s_{3})+0.1(s_{1}^{3}+s_{2}^{3}+s_{3}^{3})+\sigma\|s\|^{4}.$
Using SoSTOOLS, we verify that $m_{\mathcal{E},2}(s)-m_{\mathcal{E},2}(s^{*})$ is SoS for the tested values of $\sigma\gtrsim 10$, not not SoS for all $\sigma>0$. This observation also verifies Theorem 2.3 and Theorem 3.3. Without the cubic term, $m_{\mathcal{E}}(s)$ has a QQR structure and is therefore SoS-exact with zero layer of relaxation for all $\sigma>0$ (Theorem 3.3). After adding a small cubic perturbation, the polynomial $m_{\mathcal{E},2}(s)$ is no longer a quadratic–quartically regularized model but instead becomes a cubic–quartically regularized polynomial. In this case, the bound for $\sigma$ follows the requirements of Theorem 2.3: $\sigma$ must be sufficiently large to ensure that $B(s)$ is PSD and thus guarantee an SoS representation. Numerically, we observe that the threshold occurs around $\sigma\gtrsim 10$.
Remark 4.1.
(Non-closedness of the SoS cone under coefficient perturbations) In general, an arbitrarily small perturbation in the coefficients can move a polynomial from being SoS to non-SoS, or vice versa. This is because the cone of nonnegative polynomials is closed, while the cone of SoS polynomials is not closed in the coefficient topology [3]. Consider the well-known nonnegative but not SoS Motzkin polynomial, $M(s_{1},s_{2})=s_{1}^{4}s_{2}^{2}+s_{1}^{2}s_{2}^{4}-3s_{1}^{2}s_{2}^{2}s_{3}^{2}+s_{3}^{6}.$ Define a perturbed version $M_{\xi}(x,y)=M(x,y)+\xi(x^{6}+y^{6})$ with $\xi>0.$ For every $\xi>0$, the polynomial $M_{\xi}(x,y)$ is SoS. However, in the limit as $\xi\to 0$, $\lim_{\xi\to 0}M_{\xi}(x,y)=M(x,y),$ and $M(x,y)$ itself is not SoS.
4.2 Cubic and Quartic Maximization over the Euclidean Ball
Nesterov [46] proved that the following cubic or quartic polynomial optimization problems are NP-hard to solve globally.
Theorem 4.1 ([46, 39]).
The following polynomial optimization problems are NP-hard. Let $A_{j}\in\mathbb{R}^{n\times n}$ be positive semidefinite matrices.
  1. 1.
    Quartic maximization over the Euclidean ball:
    $\max_{s\in\mathbb{R}^{n}}\bigg\{f_{\mathcal{N}}(s):=\sum_{j=1}^{k}\left(A_{j}[s]^{2}\right)^{2}\;:\;\|s\|=1\bigg\}.$ (52)
  2. 2.
    Cubic maximization over the Euclidean ball: where $g_{3}(s)$ is a homogeneous cubic polynomial,
    $\max_{s\in\mathbb{R}^{n}}\left\{g_{3}(s)\;:\;\|s\|=1\right\}.$ (53)
  3. 3.
    Quartic minimization under quadratic constraints: where $g_{4}(s)$ is a homogeneous quartic polynomial,
    $\min_{s\in\mathbb{R}^{n}}\left\{g_{4}(s)\;:\;A_{j}[s]^{2}\geq 1,\quad j=1,\dots,k\right\}.$ (54)
    Note that this formulation is an alternative representation of (52).
We first study the quartic problem (52). Since $f_{\mathcal{N}}$ is 4-homogeneous, maximizing over the unit ball or over the sphere is equivalent: $\max_{\|s\|\leq 1}f_{\mathcal{N}}(s)=\max_{\|s\|=1}f_{\mathcal{N}}(s).$ A Lagrangian reformulation is given by
$\mathcal{L}_{1}(s,\lambda)=\sum_{j=1}^{m}\left(A_{j}[s]^{2}\right)^{2}+\lambda(\|s\|^{2}-1).$ (55)
The quartic polynomial $\mathcal{L}(s,\lambda)$ is bounded below with respect to $s$, but not with respect to $\lambda$. The Lagrangian formulation (55) is fundamentally different from our quartically regularized polynomial model in the following ways.
Remark 4.2.
Several key distinctions explain why quartic maximization over the Euclidean ball is NP-hard, whereas sufficiently quartically regularized minimization is tractable.
  1. 1.
    Unboundedness in the Lagrange multiplier. The variable pair is $(s,\lambda)$. While $\mathcal{L}(s,\lambda)$ is quartic in $s$, it contains no regularization term in $\lambda$, hence $\mathcal{L}(s,\lambda)$ is unbounded in $\lambda$.
  2. 2.
    No freedom to choose a regularization parameter. In quartic regularization, the coefficient $\sigma>0$ is a user-chosen stabilizing parameter. In contrast, here $\lambda^{*}$ is an optimization variable determined by the KKT system, not a tunable regularization parameter.
  3. 3.
    Constraint active at global solutions. Any KKT point $(s^{*},\lambda^{*})$ corresponding to a global maximizer of (55) must satisfy the boundary constraint $\|s^{*}\|=1$, implying $\lambda^{*}\neq 0$. Thus, the constraint term $\lambda(\|s\|^{2}-1)$ cannot be ignored in the global analysis.
  4. 4.
    Degeneracy when $\|g\|=0$. Consider the Lagrangian as a function of the joint variable $(s,\lambda)$. Apart from the quartic terms in $s$, $\mathcal{L}(s,\lambda)$ contains only a homogeneous cubic term and a constant term, and no first order coefficient $\|g\|=0$. As discussed in Remark 2.4, the case where the gradient vanishes is structurally special and requires careful, case-by-case analysis.
Similarly, for the cubic maximization over the Euclidean ball in (53), we can write the corresponding Lagrange multiplier formulation
$\mathcal{L}_{2}(s,\lambda)=g_{3}(s)+\lambda(\|s\|^{2}-1).$ (56)
By writing the complementarity condition in quartic form, $\mathcal{L}(s,\lambda)$ becomes a quartic polynomial in $s$. However, the four key differences highlighted in Remark 4.2 also apply here: $\mathcal{L}(s,\lambda)$ is unbounded in $\lambda$ (if $\lambda\rightarrow$-$\infty$ and $\|s\|<1$, then $\mathcal{L}(s,\lambda)\rightarrow$-$\infty$); there is no freedom to choose a regularization parameter (since $\lambda^{*}$ is an optimization variable, not a regularization weight); and the constraint is always active at a global maximizer because a cubic polynomial is unbounded; lastly, besides regularization, we have a homogeneous cubic term with $\|g\|=0$. However, if we modify (56) as
$\min_{s\in\mathbb{R}^{n}}\mathcal{L}_{c,2}(s)=g_{3}(s)+\sigma(\|s\|^{4}-1),$ (57)
where $\sigma>0$ is a regularization parameter that can be chosen a priori, then $\mathcal{L}_{c,1}(s)$ becomes a quartically regularized polynomial and is bounded below. In Figure 1, we plot the behaviour of the global minimizer $\|s^{*}\|$ as a function of the regularization parameter $\sigma$. For $\sigma$ increasing from $0.1$ to $100$, we test whether $\mathcal{L}_{c}(s)-\mathcal{L}_{c}(s^{*})$ is a sum-of-squares using GloptiPoly/MSDP. Non-SoS cases are marked with red crosses, while SoS cases are plotted by blue circles with the corresponding norm $\|s^{*}\|$ against the regularization weight $\sigma$. The right panel also includes the curve $O(\sigma^{-1/3})$ for comparison which is the theoretical bound of $\|s^{*}\|$ given in Theorem 2.2. We observe that $\mathcal{L}_{c,2}(s)-\gamma$ becomes SoS exact with zero layer of relaxation as $\sigma$ increases. The verifies our theoretical result in Theorem 2.3 that sufficiently regularized nonnegative quartic polynomials are SoS.
Figure 1: Left: Homogeneous cubic model $g_{3}(s)+\sigma\|s\|^{4}$ with random negative coefficients uniformly generated in [0,100][0,-100] and zero gradient term. Right: Non-homogeneous model $g(s)=g^{T}s+g_{3}(s)+\sigma\|s\|^{4}$, where both the cubic and linear coefficients are randomly generated in [0,100][0,-100].
Remark 4.3.
Note that assume $A_{j}\succ 0$, a similar modification can be written for (55) with
$\min_{{\tilde{s}}\in\mathbb{R}^{n}}\mathcal{L}_{1,c}({\tilde{s}},\lambda)=\sigma\|{\tilde{s}}\|^{4}+({\tilde{s}}^{T}A_{j}^{-1}{\tilde{s}}-1).$ (58)
where we use the change of variables ${\tilde{s}}^{T}{\tilde{s}}=A_{j}[s]^{2}$. We observe same numerical results with $\mathcal{L}_{1,c}$.
We would like to highlight that (57) and (56) are not equivalent and do not correspond to the same optimization problem, nor are (58) and (55). The formulation in (56) is the Lagrange multiplier expression of the constrained cubic maximization over the Euclidean ball, for which any optimal solution necessarily satisfies $\|s^{*}\|=1$. In contrast, (57) defines an unconstrained quartically regularized problem in $s$, with no restriction on the step size of its global minimizer. The subtle change in formulation alters the SoS representability and tractability of the resulting problem. This observation is consistent with known results in polynomial optimization: adding a constraint set can change the SoS representability of a problem. In the unconstrained case, many nonnegative polynomials are not SoS (as shown by Blekherman [11]), whereas Lasserre formulations often recover SoS certificates on compact domains [33]. Thus, different globalization mechanisms can indeed have a fundamental impact on the tractability of polynomial optimization problems.
4.3 Preliminary Numerical on MGH Test Set
In [75], Cartis and Zhu introduced an iterative algorithm based on SoS convex Taylor models following the work by [1]—the AR3 + SoS Convex method [75, Algorithm 1]—and established global convergence guarantees within the adaptive regularization framework. At each iteration of the AR3 + SoS Convex method, if the subproblem is not locally convex (i.e., $H\prec\delta I_{n}$), an additional quadratic regularization term, $\big(|\lambda_{\min}(H)|+\delta\big)\|s\|^{2}$, is introduced. The quartic regularization parameter, $\sigma$, is also adjusted adaptively so that the subproblem becomes SoS-convex, and at the same time, minimizing such a subproblem guarantees sufficient function value decrease. In locally convex regions, AR3 + SoS Convex method achieves the optimal function-value decrease for $p$th order method, i,e, $O(\epsilon_{g}^{1/p})$ with $p=3,5$. A limitation, however, is that the method requires the subproblem to be strongly convex at every iteration. In locally nonconvex regions, the guaranteed decrease deteriorates to $O(\epsilon_{g}^{2})$.
Motivated by our new result that sufficiently regularized nonnegative nonconvex quartic polynomials are also SoS, we propose an alternative SoS-based AR3 algorithmic framework. The resulting AR3 + SoS algorithm follows a similar adaptive update of $\sigma$ as in [75, Algorithm 1]: at each iteration $\sigma$ is increased until the AR3 subproblem becomes an SoS polynomial. Assuming uniqueness of the global minimizer, the exact solution of (Quartically Regularized Cubic Model) can then be recovered via SDP in polynomial time.
The practical variant of this algorithmic framework is given in Algorithm 1. Compared to the AR3 + SoS Convex method—where an additional quadratic regularization term must be introduced to make a nonconvex $m_{3}$. AR3 + SoS—our approach directly minimizes the original third-order Taylor model plus a quartic regularization term. In AR3 + SoS Convex, the subproblem may deviate from the true Taylor model once the quadratic regularization is added, causing the model to lose fidelity to the second- and third-order structure of $f$. In contrast, the proposed AR3 + SoS framework preserves the full third-order Taylor model, providing a more accurate local approximation in nonconvex regions. Our goal in this section is to provide preliminary evidence that sufficiently regularized AR3 subproblems can be reformulated as SoS polynomials and solved globally via SDP without introducing an additional quadratic regularization term. We emphasize that it is used only for numerical illustration; the global convergence of this variant, as well as the uniqueness of the subproblem minimizer, remains an open question.
Initialization: Choose $x_{0}\in\mathbb{R}^{n}$, Set $\sigma_{0},\sigma_{\min}>0$, $\gamma_{2}>1>\gamma_{1}>0$ , $\rho_{2}>\rho_{1}>0$.
Input: $x_{0}$, $f_{0}:=f(x_{0}),g:=\nabla_{x}f(x_{0}),H:=\nabla_{x}^{2}f(x_{0}),T:=\nabla_{x}^{3}f(x_{0})$, $\sigma_{k}:=\sigma_{0}$. Tolerance $\epsilon_{g}$ s.t $0<\epsilon_{g}\ll 1$, max iteration $k_{\max}$, $k:=0$.
Step 1: Test for termination. If $\|\nabla_{x}f(x_{k})\|\leq\epsilon_{g}$ or $k\geq k_{\max}$, terminate and output $s_{*}:=s_{k}$.
Step 2: Step computation. Compute $s_{k}=\operatorname*{argmin}_{s\in\mathbb{R}^{n}}m_{3}(s)$ by moment SDP solver where $m_{3}$ defined in (Quartically Regularized Cubic Model) with coefficients, $f_{0}=f(x_{k}),$ $g=\nabla_{x}f(x_{k}),$ $H=\nabla_{x}^{2}f(x_{k}),$ $T=\nabla_{x}^{3}f(x_{k}),$ $\sigma=\sigma_{k}$.
if Moment SDP solved successfully (status==1) then
 Extract one global optimal solution, $s_{k}$.
else
 Increase regularization: $\sigma_{k}$ by $\sigma_{k}:=\gamma_{1}\sigma_{k}$ and repeat Step 2.
Step 3: Acceptance of trial point. Compute $T_{3,x_{k}}(s_{k})$ and define $\rho_{k}=\frac{f(x_{k})-f(x_{k}+s_{k})}{f(x_{k})-T_{3,x_{k}}(s_{k})}.$
Step 4: Regularization parameter update.
if $\rho_{k}>\rho_{1}$ then
  Successful Iter. $x_{k+1}:=x_{k}+s_{k}$;
else if $\rho_{k}>\rho_{2}$ then
  Very Successful Iter. Decrease regularization: $x_{k+1}:=x_{k}+s_{k}$; $\sigma_{k+1}:=\max\{\gamma_{2}\sigma_{k},\ \sigma_{\min}\}$;
else if $\rho_{k}\leq\rho_{1}$ then
  Unsuccessful Iter. $x_{k+1}:=x_{k}$; Increase regularization: $\sigma_{k+1}:=\gamma_{2}\sigma_{k}$;
Repeat with $k:=k+1$.
Algorithm 1 Adaptive Third Order Regularization Method with SoS Taylor Models (AR3+SoS)
We present preliminary numerical results on classical unconstrained minimization problems from the Moré, Garbow, and Hillstrom (MGH) test set [42] and on other well-known examples such as Nesterov-Chebyshev. All problem initializations, as well as the first-, second-, and third-order derivatives, were generated following the specification in [8] [^6] The original implementations were in Fortran; our tests use a Matlab version obtained via a Fortran–Matlab converter [20].. All experiments were performed on an AMD Ryzen 5 7600 6-Core Processor (3.80 GHz) with 16 GB of RAM. We compare the performance of the following algorithms:
  • ARC: Adaptive Regularization with Cubics [18].
  • AR3 + ARC: Third-order method with ARC as the inner solver [19].
  • AR3 + SoS Convex: AR3 with an SoS-convex subproblem solver [75, Alg. 1].
  • AR3 + SoS: AR3 with an SoS-based Taylor subproblem solver (Algorithm 1).
Throughout this section, unless otherwise stated, the parameters for all methods were set to $\rho_{1}=0.1$, $\rho_{2}=0.9$, $\eta=0.1$ and $\gamma_{1}=0.5$, and $\gamma_{2}=2$, $\delta=0.1$, $\sigma_{0}=100$. The inner solvers used for each method were mcm from [17, Algorithm 6.1] for ARC, the ARC algorithm itself for AR3 + ARC, and the MSDP/MSOL moment-SDP solver from GloptiPoly 3 [29] with SEDUMI for AR3 + SoS Convex and AR3 + SoS. The outer-loop termination tolerance was set to $\|\nabla f\|<10^{-3}$, and the maximum iteration count at 3000. If an inner iteration was used, its tolerance was fixed at $10^{-6}$ and the maximum inner iteration count, at 3000. Appendix C.1 provides the numerical details for each problem. In performance tables, eval. refers to derivative evaluations (successful iterations only) and iter. denotes the total number of function evaluations (successful and unsuccessful). Performance profiles were used to compare the algorithms across the test set, following the framework of [25, 20, 26]. We adopt the notation of [7]: $\Gamma_{\hat{m}}(\tau_{m})$ is the proportion of problems for which method $\hat{m}$ attained an $\epsilon_{g}$-approximate solution within a computational cost no more than $\tau_{m}$ times the cost of the most efficient method, where the cost is measured in iteration or evaluation counts [^7] Some comments regarding CPU time are given at the end of the section.. Formally, if $t_{\hat{m}\hat{j}}(\epsilon_{g})$ denotes the cost incurred by method $\hat{m}$ on problem $\hat{j}$ to reach an $\epsilon_{g}$-approximate solution, we define $t^{\hat{j}}_{\min}(\epsilon_{g}):=\min_{\hat{m}}t_{\hat{m}\hat{j}}(\epsilon_{g}).$ where $\hat{m}\in\{\texttt{ARC},\texttt{AR3+ARC},\texttt{AR3+SoS Convex},\texttt{AR3+SoS}\}$}. If a method fails to satisfy the accuracy requirement within the iteration limit, we set $t_{\hat{m}\hat{j}}(\epsilon_{g})=+\infty$.
Figure 2: Performance profile plots for the three methods (ARC, AR3+ARC and AR3-SoS convex).
All test problems have dimensions ranging from $n=2$ to $n=11$, and in this range, the moment-SDP solver can cope with all subproblem dimensions. Across the test set, all methods consistently converged to the same minimizer whenever convergence happened, and each computed solution satisfied second-order necessary conditions. As can be seen in Table 2, on average [^8] These average counts exclude problems marked with * in Table 2 for the respective algorithms. AR3 + SoS required $480$ total iterations and $419$ derivative evaluations, which is less than ARC ($631$ and $628$) and comparable to AR3 + ARC ($455$ and $443$) and AR3 + SoS Convex ($494$ and $431$).
A natural question is whether the regularization parameter $\sigma$ must become extremely large in practice for the shifted AR3 subproblem to be SoS. Our numerical results indicate that this is generally not the case. Figure 3, we see that across the MGH test set, the values of $\sigma$ used by AR3 + SoS remain within a numerically stable range—typically between $10^{1}$ and $10^{6}$. These values are comparable to those produced by AR3 + SoS Convex and also broadly consistent with AR3 + ARC, except for Problem 4, which we will discuss in later paragraphs.
Figure 3: Histograms of the maximum regularization parameter $\sigma$ attained for the three methods: AR3 + SoS Convex (left), AR3 + SoS (middle), and AR3 + ARC (right). The count on the vertical axis represents the number of MGH test problems whose maximum $\sigma$ falls within each logarithmic scale bin. For most problems, the max. regularization remains within a numerically stable range ($10^{1}\sim 10^{6}$). Only a few outlier problems—notably the badly scaled cases.
Across almost all test problems, AR3 + SoS requires either the same or slightly fewer iterations and derivative evaluations than AR3 + SoS Convex. This outcome is expected: AR3 + SoS minimizes the regularized third-order Taylor model and therefore captures the local structure more accurately, whereas AR3 + SoS Convex adds a quadratic regularization term in nonconvex regions to enforce SoS-convexity, causing the model to deviate from the true Taylor approximation. Nevertheless, several MGH problems are either convex or only mildly nonconvex, and therefore the performance of the two SoS-based methods is similar on these instances. For many problems (e.g., Problems 6, 8, 9, 11, 13), the iterates remain entirely within a locally convex region. For others (e.g., Problems 5, 7, 12, 18), only the initial iterate is locally nonconvex, after which all subsequent iterates become locally convex. In these problems, the performance of AR3 + SoS and AR3 + SoS Convex is almost identical. However, a particularly interesting case is Problem 15, where the Hessian is nearly singular at every iteration. In this setting, AR3 + SoS Convex requires roughly 200 iterations and reaches a maximum regularization value of $\sigma\approx 9000$. Also, in this example, the magnitude of the regularization parameter in the iterates of AR3 + SoS Convex is highly sensitive to $\delta$. This behaviour is theoretically supported: as indicated in (34), the bound required on $\sigma$ to ensure that the subproblem in the AR3 + SoS Convex method becomes SoS-convex explicitly depends on $\delta$. In contrast, AR3 + SoS converges in only 20 iterations with a much smaller maximum regularization of $\sigma\approx 90$.
Problem 4 (Brown’s badly scaled problem) has the minimizer at $10^{4},\,1]^{T}$, and the problem is extremely ill-scaled. For ARC and AR3 + ARC, the algorithm typically proceeds with $\sigma<50$. For AR3 + SoS and AR3 + SoS Convex, as given in (29), the SoS threshold for $\sigma$ depends on $\|g\|$, $\|H\|$, $\|T\|$, which in this problem contains entries of magnitude $10^{4}$. Consequently, the SoS-enforcing $\sigma$ lies between $10^{4}$ and $10^{6}$. The algorithms remain numerically stable under this range of $\sigma$ and continue to make progress, although with smaller step sizes toward $10^{4},1]^{T}$. A simple change-of-variables scaling (see [76, 21]) would substantially reduce the required regularization.
Problem 10 and Problem 20 did not converge within the 3000 iteration limit for any of the methods. Problem 10 is known to be severely ill-conditioned, consistent with previously reported numerical difficulties [7, 20]. In Problem 20, all methods, not only the SoS-based ones, generate shrinking steps. When the 3000-iteration limit is reached, all solvers satisfy $\|\nabla f(x_{k})\|\approx 10^{-2}$ and are clearly converging but have not yet reached high accuracy. Increasing the maximum iteration limit to 6000 allows AR3 + ARC and AR3 + SoS Convex to converge fully, but ARC still fails to satisfy the gradient tolerance. Thus, the stagnation is intrinsic to the problem geometry rather than a consequence of the SoS requirement or large $\sigma$.
4.3.1 Two illustrative examples
Nesterov-Chebyshev.
This function was introduced by Nesterov [31, 48] and it is defined by
$f_{r}(x)=0.25(x_{1}-1)^{2}+\sum_{i=1}^{n-1}(x_{i+1}-2x_{i}^{2}+1)^{2},$
a fourth-degree polynomial with a global minimizer $x_{\text{glob}}^{*}=(1,\ldots,1)^{T}$ with $f_{r}(x^{*})=0$. At any other $x\in\mathbb{R}^{n}$, the gradient $\nabla f_{r}(x)$ is nonzero. Although the function has a unique stationary point (i.e., the global minimizer), the function contains many near-stationary regions with very small gradients, leading to slow convergence for first- and second-order methods. For $n=2$, the Chebyshev–Rosenbrook function reduces to the form of Rosenbrock function [^9] Note that Problem 1 in the MGH test set corresponds to the Rosenbrock function with different coefficients scaling, which is $(x_{1}-1)^{2}+100(x_{2}-x_{1}^{2})^{2}$.. We apply the AR3 + SoS method to this function. Figure 4(a) shows a heat map of convergence behaviour over a fine grid of initializations in $4,4]^{2}$, where the color denotes the distance between the final iterate and the global minimizer. Although each subproblem is solved globally via the moment-SDP solver, global convergence of the outer iteration is not guaranteed. Nonetheless, Figure 4(a) shows that, for all initial points in the grid, the AR3 + SoS method consistently converges to $x_{\text{glob}}^{*}$, typically within fewer than $30$ outer iterations. These preliminary results suggest that globally solving the subproblem can help steer the outer iteration toward the global minimizer.
(a) Basins of attraction for the Chebyshev–Rosenbrock.
(b) Basins of attraction for the Beale function
Figure 4: Global convergence behaviours across different initializations. Dark regions indicate convergence to the global minimizer, while lighter regions indicate no convergence after reaching the maximum number of iterations.
Solving the Chebyshev–Rosenbrock function for higher values of $n$ is well known to be challenging, as iterates can easily become trapped in flat regions where the gradients are very small. Classical methods such as Newton’s method and BFGS often require a large number of iterations when attempting to reach high accuracy solutions [27]. We delegate to future work the investigation of the performance of the AR3+SoS method on this example in higher-dimensional settings.
The Beale function.
The Beale function, which is the fifth problem in the MGH test set [42],
$f_{b}(x_{1},x_{2})=(1.5-x_{1}+x_{1}x_{2})^{2}+(2.25-x_{1}+x_{1}x_{2}^{2})^{2}+(2.625-x_{1}+x_{1}x_{2}^{3})^{2}$
is a standard benchmark for testing robustness to initialization. Although it has a unique global minimizer $\tilde{x}_{\mathrm{glob}}^{*}=(3,0.5)^{T}$ and no other local minima, its landscape is highly nonconvex, with long, narrow valleys and large flat regions that typically stall first- and second-order methods. Following the setup of [1, Sec. 5.2], we evaluate global convergence behaviour over a uniform grid on $4,4]^{2}$, running each method for at most $350$ iterations. In Figure 4(b), dark regions correspond to initializations that converge (approximately) to the global minimizer. Lighter regions represent runs that fail to reach the minimizer and terminate only because the iteration budget is exhausted; in these cases, the distance to the minimizer remains $\mathcal{O}(1)$. Increasing the iteration cap from 3,000 to 20,000 iterations does not improve these non-convergent cases, as the iterates remain trapped in flat regions—consistent with the behaviour reported in [1, Sec. 5.2]. As shown in Figure 5, the region of initializations that converge to the global minimum (yellow area) AR3+SoS and AR3+SoS Convex are substantially larger than those of Newton method [^10] For Newton and the third-order Newton method, full numerical results appear in [1, Sec. 5.2]., indicating an expansion of convergence regions when moving from second-order to third-order methods.
(a) Newton [1, Fig. 5]
(b) 3rd-order Newton [1]
(c) AR3+SoS Convex
(d) AR3+SoS
Figure 5: Global convergence behaviours of different initialization for minimizing the Beale function (a) the Newton method, (b) the third-order Newton method, and (c) AR3 + SoS Convex with $\delta=0.01$. Yellow indicates $\|x^{k}-\tilde{x}_{\mathrm{glob}}^{*}\|/\|\tilde{x}_{\mathrm{glob}}^{*}\|\leq 10^{-3}$ where $x^{k}$ is the outcome after 350 iterations. Blue indicates all other behaviors, including divergence or convergence to non-optimal points
4.4 Sensor Network Localization Problem
Quartic optimization problems arise in various engineering applications such as independent component analysis [14], blind channel equalization in digital communication [40], and sensor network localization [10]. We consider the sensor localization problem, where $\cal A$} and $\cal S$} denote the sets of anchor nodes and sensor nodes, with cardinalities $|{\cal A}|$ and $|{\cal S}|$, respectively. The goal is to estimate the sensor positions $x_{i}\in\mathbb{R}^{3}$ for $i\in{\cal S}$ by minimizing the quartic objective
$\min_{x_{i}\in\mathbb{R}^{3},\,i\in{\cal S}}f_{{\cal S}}(x):=\sum_{i,j\in{\cal S}}\big(\|x_{i}-x_{j}\|^{2}-d_{ij}^{2}\big)^{2}+\sum_{i\in{\cal S},\,j\in{\cal A}}\big(\|x_{i}-s_{j}\|^{2}-d_{ij}^{2}\big)^{2},$ (59)
where $d_{ij}$ are (possibly noisy) distance measurements between nodes, $s_{j}$ are the known positions of the anchors, and $x=[x_{1}^{T},x_{2}^{T},\dotsc,x_{|{\cal S}|}^{T}]^{T}\in\mathbb{R}^{3|{\cal S}|}.$ By homogenizing (59), one obtains a quartic minimization problem with a single quadratic constraint, fitting the structure of (54). As a resuult, (59) is known to be NP-hard [39] [^11] Indeed, even deciding whether the optimal value of (59) is zero is NP-hard.. The relative sizes of the sensor and anchor sets strongly influence the difficulty of the localization problem. When anchors outnumber sensors ($|{\cal A}|>|{\cal S}|$), the system is overdetermined and relatively easier. In practical settings, however, only a few anchors are available (e.g., GPS-equipped nodes), and the common regime is $|{\cal A}|\leq|{\cal S}|$. This underdetermined case is significantly more challenging: the resulting optimization landscape is highly nonconvex, often sensitive to measurement noise, and typically contains many spurious local minima. Moreover, due to geometric symmetries, multiple global minimizers may exist. For example, even in the simplest case with two anchors and two sensors, the problem admits several symmetric global solutions, and the rank condition required for moment-SDP solvers (e.g., GloptiPoly/MSDP) fails—preventing extraction of the global minimizer from the relaxation.
The quartic sensor–localization model is well suited for third-order methods. Although the problem is highly nonconvex, its third-order tensor has a structured (and often low-rank) form determined by the anchor geometry. Moreover, the third-order tensor can be applied to directions without explicitly forming the full dense tensor; see Section C.3 for details. We test four methods on this problem: AR3 + SoS, fminunc, ARC, and AR3 + ARC.
We try on the smallest “hard” case, with $|{\cal S}|=|{\cal A}|=2$ a various level of noise $10^{-5}-10^{0}$. In this setting, $f_{{\cal S}}$ is a quartic polynomial in $n=2\times 3=6$ variables. Our experiments show that, from the same initialization, AR3+SoS usually reduces the objective value the most giving near zero function value (below $O(10^{-5})$ for noise level $10^{-5}-10^{-1}$). In contrast, ARC, AR3 + ARC, and fminunc typically converge to a higher objective function value, which correspond to a local minimizers. It is worth noting that, we also tried reformulate the 66–variable quartic directly (59) in SoS program and solve it with GloptiPoly/MSDP. However, at $|{\cal S}|=|{\cal A}|=2$, the localization problem already have too much global minimizers, so the moment matrix in the Lasserre relaxation fails to satisfy the flat extension (rank) condition and a global minimizer cannot be extracted from the SDP solution. MSDP returning status ==0 which means SoS relaxation solved correctly, but cannot recover the global minimizer since the rank condition fails. In contrast, the AR3+SoS algorithm converges numerically to an approximate global minimizer.
Trule Noise Level $10^{-6}$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $^{-2} 1001$$^{0}$
AR3 + SoS 11/11 11/11 18/17 19/19 20/19 24/22
fminunc 8/ 6 12/11 9/ 7 18/10 15/11 13/ 7
ARC 17/12 27/19 20/17 19/16 26/18 8/ 7
AR3 + ARC 9/ 7 10/ 7 13/ 9 10/ 7 14/10 28/16
Figure 6: Comparison of four methods on the sensor–localization problem with 22 anchors and 22 sensors ($n=6$ variables). Distances $d_{i}j$ are perturbed by adding Gaussian noise of magnitude $10^{-6}$ to $10^{0}$. All algorithms start from the anchor initialization. Left: Function values at $x^{*}$ for different noise levels. Right: Iteration / evaluation counts for each method at the same noise levels.
Scalability and CPU time considerations.
Our experiments are limited to relatively small-scale problems because the inner subproblems are solved using moment-SDP relaxations, which are practical only for small polynomial systems ($n<10$). In comparison with other methods, AR3 + SoS is more expensive in terms of CPU time per iteration. Although SoS relaxations offer a principled route to global optimality, the SDP size $N=\binom{n+4}{4}$ limits their applicability. In settings where the polynomial subproblems admit structured sparsity, a sparsity-adapted hierarchy of relaxations can significantly reduce SDP size [13, 37, 41, 69]. Additionally, several scalable alternatives to full SoS relaxations have been proposed, such as (scaled) diagonally dominant SoS (DSoS/SDSoS) and bounded-degree SoS (BSoS) [3, 38, 69, 70, 72]. Replacing the SDP solves in Algorithm 1 with more scalable LP- or SOCP-based approaches or exploiting problem structure to generate smaller SDPs is a promising direction for making the approach applicable to larger-scale problems [3]. The focus of this paper is on establishing a theoretical framework that enables a higher-order Newton-type algorithm to achieve global convergence guarantees while maintaining polynomial-time complexity per iteration. In particular, our work—together with the recent results of Ahmadi [1]—provides one of the first guarantees that a third-order regularization scheme can be paired with a subproblem that is globally solvable in polynomial time. The development of scalable variants based on approximate SoS relaxations or structured sparsity is left for future work.
5 Conclusion
In this paper, we studied the interplay between regularization, convexification, and Sum-of-Squares (SoS) representability for quartically regularized cubic polynomials arising from high-order adaptive regularization methods. We showed that, under sufficiently strong regularization, a nonconvex nonnegative quartically regularized polynomials admits an SoS representation. This establishes, for the first time, explicit SoS certification results for a family of nonconvex quartically regularized models.
In addition, we identified several structured subclasses of quartically regularized cubic polynomials for which global optimality of the model implies that its nonnegativity is certified by a Sum-of-Squares (SoS) decomposition, for all values of the regularization parameter $\sigma$. In this sense, our results extend Hilbert’s classical characterizations of nonnegative polynomials to a new subfamily motivated by high-order tensor methods. We further propose an algorithmic framework that combines SoS-based Taylor models with adaptive regularization techniques.
Several directions for future research naturally arise from this work. These include establishing global complexity guarantees for the AR33+SoS algorithmic framework, and investigating tight semidefinite relaxations with reduced dimensionality. In particular, recent work by [73] leverages global sufficient optimality conditions for cubic quartically regularized (CQR) polynomials [21] to derive a tight SDP formulation of size $n+1$. Another promising direction is to exploit additional structural properties to develop more scalable variants using alternatives [38, 69, 70, 72].
Acknowledgments: This work was supported by the Hong Kong Innovation and Technology Commission (InnoHK Project CIMDA). Coralia Cartis’s research was supported by the EPSRC grant EP/Y028872/1, Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI.
Appendix A Supplementary details for the main results
A.1 Comparison of Tensor Norms
For a symmetric third-order tensor $T\in\mathbb{R}^{n\times n\times n}$, define its cubic (spectral) norm as
$\|T\|:=\max_{\|s\|=1}|T[s]^{3}|=\max_{\|s\|=1}|\langle T,\,s\otimes s\otimes s\rangle|.$
By the Cauchy–Schwarz inequality for tensor inner products, we have
$|T[s]^{3}|=|\langle T,\,s\otimes s\otimes s\rangle|\leq\|T\|_{F}\,\|s\otimes s\otimes s\|_{F}=\|T\|_{F}\,\|s\|^{3}=\|T\|_{F}.$
Hence,
$\boxed{\|T\|\leq\|T\|_{F}.}$
Moreover, using the mode-1 unfolding $T_{(1)}\in\mathbb{R}^{n\times n^{2}}$, one obtains
$T[s]^{3}=s^{T}T_{(1)}(s\otimes s)\quad\Rightarrow\quad|T[s]^{3}|\leq\|T_{(1)}\|_{2},$
and therefore
$\|T\|\leq\min\{\|T_{(1)}\|_{2},\,\|T_{(2)}\|_{2},\,\|T_{(3)}\|_{2}\}\leq\|T\|_{F}.$
Equality $\|T\|=\|T\|_{F}$ holds if and only if $T$ is (up to a sign) a rank-one symmetric tensor of the form $T=\alpha\,u\otimes u\otimes u$ with $\|u\|=1$. A simple lower bound follows from the diagonal entries:
$\max_{1\leq i\leq n}|T_{iii}|\leq\|T\|.$
Thus, $\|T\|$ is always upper bounded by the Frobenius norm, attains equality for rank-one symmetric tensors, and provides a tighter measure of the dominant cubic interaction captured by $T$.
A.2 A length of a local minimizer $s^{*}$ is a non-increasing function of $\sigma$
Theorem A.1.
Suppose that $s^{*}=s^{*}(\sigma)\neq 0$ is any second-order local minimizer of the model $m_{3}(s)$. Then the length of the minimizer $v(\sigma)=\|s^{*}(\sigma)\|$ is a non-increasing function of $\sigma$.
Proof.
Since $s^{*}(\sigma)\neq 0$ a second-order local minimizer of the model $m_{3}(s)$, then
$\nabla m_{3}\big(s^{*}(\sigma)\big)=0,\qquad Hs^{*}(\sigma)+\frac{1}{2}T[s^{*}(\sigma)]^{2}+\sigma\|s^{*}(\sigma)\|^{2}s^{*}(\sigma)=-g.$ (60)
Differentiating (60) with respect to $\sigma$ gives
$H\nabla_{\sigma}s^{*}(\sigma)+T[s^{*}(\sigma)]\nabla_{\sigma}s^{*}(\sigma)+\sigma\|s^{*}(\sigma)\|\nabla_{\sigma}s^{*}(\sigma)+2s^{*}(\sigma)^{T}s^{*}(\sigma)\nabla_{\sigma}s^{*}(\sigma)+\|s^{*}(\sigma)\|^{2}s^{*}(\sigma)=0$ (61)
and therefore
$\nabla^{2}m_{3}(s^{*}(\sigma))\nabla_{\sigma}s^{*}(\sigma)+\|s^{*}(\sigma)\|^{2}s^{*}(\sigma)=0.$ (62)
Let $v(\sigma)=\|s^{*}(\sigma)\|$, then $v^{\prime}(\sigma)=\|s^{*}(\sigma)\|^{-1}\nabla_{\sigma}s^{*}(\sigma)^{T}s^{*}(\sigma)$. Premultiplying (62) by $\nabla_{\sigma}s^{T}(\sigma)$ and dividing by $\|s^{*}(\sigma)\|$ gives
$v^{\prime}(\sigma)=\|s^{*}(\sigma)\|^{-1}\nabla_{\sigma}s^{*}(\sigma)^{T}s^{*}(\sigma)=\frac{-\nabla_{\sigma}s^{*}(\sigma)^{T}\nabla^{2}m_{3}(s^{*}(\sigma))\nabla_{\sigma}s^{*}(\sigma)}{\|s^{*}(\sigma)\|^{2}}\leq 0.$
Since we have seen that $\nabla^{2}m_{3}(s^{*}(\sigma))$ is positive semidefinite. Thus, $s^{*}(\sigma)$ is a non-increasing function of $\sigma$. ∎
Remark A.1.
Theorem A.1 can be extended to $p$th order for any second-order local minimizer of $m_{p}(s)$.
A.3 Global Optimality Conditions for $m_{3}$
For a general $m_{3}$, global optimality conditions were established in [76]. Let $s^{*}\in\mathbb{R}^{n}$ satisfy $\nabla m_{3}(s^{*})=0$. If $s^{*}$ is a global minimum, then
$B_{N}(s^{*}):=H+\frac{2}{3}T[s^{*}]+\sigma\|s^{*}\|^{2}I_{n}+\frac{\|T\|}{3}\|s^{*}\|I_{n}\succeq 0.$ (Necessary cond.)
If the following is satisfied,
$B_{S}(s^{*}):=H+\frac{2}{3}T[s^{*}]+\sigma\|s^{*}\|^{2}I_{n}-\frac{\|T\|}{3}\|s^{*}\|I_{n}-\frac{\|T\|^{2}}{18\sigma}I_{n}\succeq 0,$ (Sufficient cond.)
then $s^{*}$ is a global minimum.
Theorem A.2.
Assume that $m_{3}$ is locally nonconvex at $s=0$, namely $\nabla m_{3}(0)=H$ is not positive definite. Let $s^{*}\in\mathbb{R}^{n}$ satisfy $\nabla m_{3}(s^{*})=0$. If $\sigma$ is chosen such that
$\sigma\geq 9\max\bigg\{\frac{27\lambda_{*}^{3}}{\|g\|^{2}},\quad\frac{\|H\|^{3}}{\|g\|^{2}},\quad\bigg(\frac{\|T\|^{3}}{\|g\|}\bigg)^{1/2}\bigg\},$ (63)
where $\lambda_{*}:=\min\{-\lambda_{\min}(H),0\}\geq 0$, $\|H\|$, $\|T\|$ are defined in (22)–(23). Then, the following statements hold:
  1. 1.
    $\big(\frac{\|g\|}{3\sigma}\big)^{1/3}\leq\|s^{*}\|\leq 2\big(\frac{\|g\|}{\sigma}\big)^{1/3}$ and $\sigma\geq 3\max\bigg\{\lambda_{*}\|s^{*}\|^{-2},\|T\|\,\|s^{*}\|^{-1}\bigg\},$
  2. 2.
    $B_{S}(s^{*})\succeq 0$, $s^{*}$ is the global minimizer of $m_{3}$, the necessary condition (Necessary cond.) and sufficient condition (Sufficient cond.) coincide.
Proof.
To prove result 1: If $\sigma$ satisfies (63), then $\sigma$ satisfies the condition in Theorem 2.2, and the bound on $\|s^{*}\|$ in the first result follows directly from Theorem 2.2. Moreover, from (63), we have
$\sigma^{1/3}\geq 3^{5/3}\|g\|^{-2/3}\lambda_{*},\qquad\sigma^{2/3}\geq 3^{4/3}\|g\|^{-1/3}\|T\|.$
We deduce that
$\sigma\geq 3\max\bigg\{3^{2/3}\sigma^{2/3}\|g\|^{-2/3}\lambda_{*},\quad 3^{1/3}\sigma^{1/3}\|g\|^{-1/3}\|T\|\bigg\}\geq 3\max\bigg\{\lambda_{*}\|s^{*}\|^{-2},\quad\|T\|\|s^{*}\|^{-1}\bigg\}$ (64)
where the last inequality uses $\|s^{*}\|\geq\big(\frac{\|g\|}{3\sigma}\big)^{1/3}$ (i.e., $3^{1/3}{\sigma}^{1/3}\|g\|^{-1/3}\geq\|s^{*}\|^{-1}$) from Theorem 2.2.
To prove result 2: Given (64), we have
$\sigma>6^{-1/2}\|T\|\|s^{*}\|^{-1}\qquad$ $\Rightarrow$ $\qquad\frac{\sigma}{3}\|s^{*}\|^{2}>\frac{{\|T\|}^{2}}{18\sigma},$
$\sigma\geq 3\|T\|\|s^{*}\|^{-1}\qquad$ $\Rightarrow$ $\qquad\frac{\sigma}{3}\|s^{*}\|^{2}\geq\|T\|\|s^{*}\|,$
$\sigma\geq 3\lambda_{*}\|s^{*}\|^{-2}\qquad$ $\Rightarrow$ $\qquad\frac{\sigma}{3}\|s^{*}\|^{2}\geq\lambda_{*}.$
We deduce that
$B_{S}(s^{*})\succeq\bigg(\lambda_{*}-\|T\|\|s^{*}\|+\sigma\|s^{*}\|-\frac{\|T\|^{2}}{18\sigma}\bigg)I_{n}\succeq 0.$
Since $\nabla m_{3}(s^{*})=0$ and $B_{S}(s^{*})\succeq 0$, it follows that $s^{*}$ is the global minimizer of $m_{3}$. Furthermore, as $s^{*}$ is a global minimizer, the necessary condition $B_{N}(s^{*})\succeq 0$ also holds. By construction, we already have $B_{N}(s^{*})\succeq B_{S}(s^{*})\succeq 0$, and thus the necessary condition (Necessary cond.) and the sufficient condition (Sufficient cond.) coincide. ∎
Appendix B Supplement for Classes of Quartically Regularized Polynomial
B.1 Proof for Theorem 3.3
Proof.
Assume $T=0$, since $s^{*}$ is the global minimum, we have $\nabla m_{3}(s^{*})=0$, according to (5), we can write
$m_{3}(s^{*}+v)-m_{3}(s^{*})=\frac{1}{2}\omega\underbrace{\begin{bmatrix}0&0&0&0&\\ 0&\nabla^{2}m_{3}(s^{*})&\sigma s\mathbf{1}^{T}&0\\ 0&\sigma\mathbf{1}s^{T}&\frac{\sigma}{2}\mathbf{1}\mathbf{1}^{T}&0\\ 0&0&0&0\end{bmatrix}}_{Q_{b}}\omega$ (65)
where $\phi_{2}$ defined as in (2), $\mathbf{1}=[1,\dotsc,1]^{T}\in\mathbb{R}^{n}$ and $\mathbf{1}\mathbf{1}^{T}\in\mathbb{R}^{n\times n}$ is a matrix with all entries equal to 1s. Let $\omega\in\mathbb{R}^{\frac{1}{2}(n+2)(n+1)}$ be any vector, we write $\omega=[1,v,u,z]^{T}$ as defined in (3). For any $\omega$, or equivalently for any $(u,v,z)$, we have
$2\omega^{T}Q_{b}(v)\omega$ $=$ $\nabla^{2}m_{3}(s^{*})[v]^{2}+2\sigma v^{T}s^{*}\sum_{i=1}^{n}u_{i}+\frac{\sigma}{2}\sum_{i=1}^{n}u_{i}^{2}+\sigma\sum_{1\leq i<j\leq n}u_{i}u_{j}$ (66)
$=$ $\nabla^{2}m_{3}(s^{*})[v]^{2}+\frac{\sigma}{2}\bigg(\sum_{i=1}^{n}u_{i}+2v^{T}s^{*}\bigg)^{2}-2\sigma(v^{T}s^{*})^{2}$
$=$ $\big[\underbrace{\nabla^{2}m_{3}(s^{*})-2\sigma s^{*}{s^{*}}^{T}}_{\succeq 0}\big][v]^{2}+\frac{\sigma}{2}\bigg(\sum_{i=1}^{n}u_{i}+2v^{T}s^{*}\bigg)^{2}\geq 0$
where the last inequality uses the second-order global optimality condition for $s^{*}$ given in Lemma 3.3 (i.e., $H+\sigma\|s^{*}\|I_{n}\succeq 0$). Thus, we obtained that $Q_{b}\succeq 0$ and therefore $m_{3}(s^{*}+v)-m_{3}(s^{*})$ is SoS. ∎
Remark B.1.
The bound in (66) is tight, as we can set $v=-s^{*}$ and $\sum u_{i}=\|s^{*}\|^{2}$. When $n=1$, Theorem 3.3 becomes the result in Remark 3.1.
Appendix C Supplement for Numerical Results
C.1 Experiment with MGH Test Set
In Table 2, * indicates that the algorithm failed to make progress after 3000 iterations. The Iter. and Eval. for AR3+SoS and AR3+SoS Convex excludes the unsolved SoS problems—these columns refer specifically to function evaluations and derivative evaluations. Problems marked with * are excluded from the average count. $\dagger$} indicates that the algorithm timed out: $\|\nabla f\|$ was reduced below $10^{-2}$ but did not reach the prescribed outer-loop tolerance of $10^{-3}$.
Table 2: Comparison of AR3+SoS Convex, AR3+SoS (Algorithm 1), ARC, and AR3+ARC across 20 test problems.
Problem Dim AR3 + SoS C AR3 + SoS ARC AR3 + ARC
$n$ Iter. Eval. Iter. Eval. Iter. Eval. Iter. Eval.
1 2 2656 2305 2656 2305 2317 2317 2354 2311
2 2 96 84 96 84 70 70 58 57
3 2 29 12 27 10 26 13 31 12
4 2 * * * * 1372 1372 32 31
5 2 67 59 67 59 57 57 58 57
6 2 113 100 113 100 104 104 119 99
7 3 815 711 792 691 779 779 717 716
8 3 825 723 825 723 707 707 726 725
9 3 19 11 18 11 24 17 17 12
10 3 * * * * * * * *
11 3 51 41 56 47 29 28 50 40
12 3 26 26 26 26 24 24 26 25
13 4 231 207 231 207 209 209 208 206
14 4 93 84 90 81 53 40 119 63
15 4 230 200 22 22 24 22 23 22
16 4 121 121 121 121 91 91 92 85
17 5 335 288 335 288 3000 2988 173 148
18 6 16 16 17 17 13 13 17 16
19 11 167 155 152 140 96 83 831 800
20 6 3000 2605 3000 2605 3000 3000 3000 3000
Average 494 431 480 419 631 628 455 443
C.2 Supplementary Results for Perturbing $m_{\mathcal{E}}$
Table 3: Comparison of Global Minima Found by MoM and $\gamma^{*}$ Found by SoS
Objective Global Mins Found by MoM $\gamma^{*}$ Found by SoS
$m_{\mathcal{E}}$ - $1-3.6655\times 10^{-10}$
$\hat{m}_{\mathcal{E},1}$ $\pm[0.7071,-0.7071,-0.0000]^{T}$ $1-4.0000\times 10^{-5}$
$\hat{m}_{\mathcal{E},2}$ - run 1 $\pm[0.7809,-0.1830,-0.5979]^{T}$ $1-6.6080\times 10^{-6}$
$\hat{m}_{\mathcal{E},2}$ - run 2 $\pm[0.2089,0.5790,-0.7879]^{T}$ $1-2.1461\times\ 10^{-5}$
$\hat{m}_{\mathcal{E},2}$ - run 3 $\pm[0.7107,-0.0072,-0.7035]^{T}$ $1-4.3092\times 10^{-5}$
$\hat{m}_{\mathcal{E},2}$ - run 4 $\pm[0.8096,-0.3135,-0.4962]^{T}$ $1-5.6846\times 10^{-5}$
$\hat{m}_{\mathcal{E},2}$ - run 5 $\pm[0.7903,-0.2176,-0.5728]^{T}$ $1-2.8288\times 10^{-5}$
C.3 Derivative and Tensor for Sensor–Localization model
Let $S$ be the set of sensor indices, $A$ the set of anchors, and let each sensor/anchor position lie in $\mathbb{R}^{3}$. For $(i,j)\in S\times S$ and $(i,j)\in S\times A$, define the residuals
$r_{ij}(x)\;=\;\|x_{i}-x_{j}\|^{2}-d_{ij}^{2},\qquad\tilde{r}_{ij}(x)\;=\;\|x_{i}-s_{j}\|^{2}-d_{ij}^{2},$
so that the objective can be written as the quartic polynomial
$f(x)=\sum_{i,j\in S}r_{ij}(x)^{2}+\sum_{i\in S,\,j\in A}\tilde{r}_{ij}(x)^{2}.$
Each term is a composition of a squared norm and a quadratic residual. Consequently, the derivatives can be obtained by simple applications of the chain rule. For example, the gradient with respect to $x_{i}$ is
$\nabla_{x_{i}}f(x)=4\sum_{j\in S}r_{ij}(x)\,(x_{i}-x_{j})+4\sum_{j\in A}\tilde{r}_{ij}(x)\,(x_{i}-s_{j}),$
and the Hessian consists of rank–one and rank–two updates built from the displacement vectors $x_{i}-x_{j}$ and $x_{i}-s_{j}$. Differentiating once more, the third derivative $\nabla^{3}f(x)$ is a structured $3|{\cal S}|\times 3|{\cal S}|\times 3|{\cal S}|$ tensor whose nonzero blocks are linear combinations of tensors of the form $(x_{i}-x_{j})\otimes(x_{i}-x_{j})\otimes(x_{i}-x_{j}),$ and $(x_{i}-s_{j})\otimes(x_{i}-s_{j})\otimes(x_{i}-s_{j}),$ together with lower–rank symmetric permutations. Thus, although the problem is highly nonconvex, its third–order tensor has a highly structured form (and potentially low rank depending on the location of the anchor). Also, the third–order tensor can be applied to directions without ever forming the full dense tensor explicitly. This makes the model particularly suitable as a testbed for third–order methods.